### v1

# VLA-Cache: Towards Efficient Vision-Language-Action Model via Adaptive Token Caching in Robotic Manipulation

Abstract
Vision-Language-Action (VLA) model can process instructions and visual perception to directly generate actions as output in an end-to-end fashion due to its strong multi-modal reasoning capabilities. While the performance of VLA models is promising, their computational cost can be substantial. This raises challenge for applying them on robotics tasks, which requires real-time decision-making to respond quickly to environmental changes. Since robotic control involves sequential decision-making, the visual input often exhibits minimal variation between successive steps. A natural idea is to reuse the computational results of unchanged visual tokens from the last step. Motivated by this idea, we propose VLA-Cache, an efficient vision-language-action model. VLA-Cache incorporates a token-selection mechanism that compares the visual input at each step with the input from the previous step, adaptively identifying visual tokens with minimal changes. The computational results for these unchanged tokens are then reused in subsequent steps via KV-cache, thereby significantly improving the efficiency of the VLA-Cache model. Experimental results on both simulation (e.g., LIBERO benchmark and SIMPLER) and real-world robot valid VLA-Cache can achieve practical acceleration with minimal sacrifice in success rate.
1 Introduction
Learning a robust and generalizable policy for robotic manipulation through policy learning has long been a challenging problem (Kim et al., 2020), with traditional reinforcement learning approaches (Choi et al., 2024; Bica et al., 2021) often suffering from poor robustness and limited generalization. Recently, the rapid advancement of foundational Vision-Language Models (VLMs) (Awadalla et al., 2023; Liu et al., 2024b) has demonstrated remarkable capabilities in multimodal understanding and generalization. Leveraging large-scale real-world robotic datasets (O’Neill et al., 2024; Fang et al., 2024), pioneering works (Zitkovich et al., 2023; Octo Model Team et al., 2024; Niu et al., 2024; Kim et al., 2024) have introduced Vision-Language-Action (VLA) models, which integrate vision and language modalities to directly generate robotic actions in an end-to-end manner. This emerging paradigm holds great promise for enhancing the adaptability and generalization of robotic control systems, but leaves a large computational demand.
Figure 1: During the inference of the VLA model, static tokens of the input image remain largely consistent across steps. This consistency allows for caching the computations of these tokens from the previous step and improving the efficiency of inference.
To mitigate the extensive cost of VLA models, existing works often adopt generic acceleration techniques, such as model lightweighting (Wen et al., 2024), quantization (Park et al., 2024), and early exit (Yue et al., 2024). However, these approaches fail to account for the unique characteristics of VLA models, leading to suboptimal efficiency gains and limited performance improvements.
In this paper, we argue that robotic manipulation, characterized by sequential action-scene interactions, inherently prioritizes visual inputs related to the robotic arm and the target object, while the background remains largely static and less critical. As illustrated in Figure 1, the static tokens of the input image remain largely consistent across steps. In this context, the background tokens, which contribute significantly to computational overhead and repeat across sequential steps in the VLA process, are often redundant.
As a result, we propose VLA-Cache, a novel approach that identifies and reuses static tokens across sequential VLA steps. Specifically, static tokens in each step refer to visual tokens that exhibit minor changes compared to the previous step. Instead of recomputing these tokens, VLA-Cache directly substitutes them with their cached representations from the preceding step, significantly reducing redundant computations and improving inference efficiency. Nevertheless, we observe that model precision is highly sensitive to task-relevant tokens (i.e., those related to the robotic arm and the target object) in the ablation study, where even a slight positional shift can lead to a significant performance drop. To address this, we introduce a fine-grained selection scheme that filters out task-relevant tokens from the static token set, ensuring that these critical tokens continue to undergo full computational interactions, thereby preserving model accuracy while maintaining efficiency.
Additionally, based on the observation that attention distributions within the VLA decoder vary significantly across different layers, we propose a layer-adaptive strategy that adjusts the fraction of reused tokens based on each layer’s attention concentration. We evaluate VLA-Cache on a diverse set of robotic manipulation tasks in the LIBERO benchmark, where it demonstrates over 1.7× acceleration compared to the baseline, with only slight drop on the success rate. Furthermore, VLA-Cache is deployed on the Kinova Jaco robot arm, showing that it can achieve practical acceleration in real-world scenarios.
2 Related Work
Vision-Language-Action Models. Large-scale vision-language models (VLMs) have significantly advanced multimodal learning by integrating image understanding and language reasoning (Liu et al., 2024b; Bai et al., 2023). Extending these capabilities, Vision-Language-Action (VLA) models (Zitkovich et al., 2023; Li et al., 2023) incorporate an action modality, enabling end-to-end visuomotor control. These models typically adopt large VLM backbones (Touvron et al., 2023) and fine-tune them on robot data (O’Neill et al., 2024), with approaches varying from discretizing actions as language-like tokens (Kim et al., 2024; Chi et al., 2023) to incorporating specialized diffusion policy heads (Black et al., 2024). Despite their effectiveness in tasks like object retrieval and assembly (Huang et al., 2024; Zhao et al., 2023), VLA models demand substantial computation, making real-time deployment challenging, particularly in resource-constrained environments.
Acceleration for Vision-Language Models. Acceleration techniques for VLMs, such as quantization (Xie et al., 2024) and pruning (Lin et al., 2024a), have shown success in vision-language tasks but often overlook the real-time demands of action generation. Token-level methods like FastV (Chen et al., 2025) and SparseVLM (Zhang et al., 2024) optimize inference by pruning or merging redundant tokens, but their effectiveness in VLA tasks, which require short and highly specialized action sequences, remains unclear. Some works address efficiency by modifying VLA architectures, such as DeeR-VLA (Yue et al., 2024), which dynamically adjusts inference depth, and QAIL (Park et al., 2024), which integrates quantization-aware training. Others, like RoboMamba (Liu et al., 2024c) and TinyVLA (Wen et al., 2024), focus on replacing traditional attention mechanisms or training compact models from scratch, often requiring re-training and additional data collection. In contrast, VLA-Cache offers a training-free acceleration approach by selectively caching static tokens and recomputing only dynamic or task-relevant ones, preserving crucial temporal cues while reducing inference latency and improving real-time robotic control efficiency.
3 Preliminary
3.1 Vision-Language-Action Models
Vision-Language-Action (VLA) models extend large-scale VLMs to generate robotic actions based on multimodal inputs, including visual observations and task instructions. These models are typically trained on large-scale datasets, such as Open X Embodiment (O’Neill et al., 2024), which contains over a million real-world robot trajectories. The training pipeline involves a pre-trained vision encoder (Liu et al., 2021; Radford et al., 2021) extracting image features, which are projected into the input space of a Large Language Model (LLM). The LLM, fine-tuned alongside a projector module, encodes these inputs to predict actions.
During inference, VLA models process sequential frames to iteratively guide the robot, but the full Transformer-based LLM inference at each step results in substantial computational overhead. The quadratic complexity of self-attention, combined with large key-value (KV) caches (Ge et al., 2023), leads to high memory consumption and slow inference speeds, posing challenges for real-time deployment. While model compression techniques such as quantization (Lin et al., 2024b) and pruning (Park et al., 2024; Yue et al., 2024) have been explored to reduce computational cost, they typically require re-training and may degrade performance, limiting their practicality in real-time applications. To mitigate these inefficiencies, our approach leverages token caching mechanisms to accelerate inference in a training-free manner, reducing computational cost while preserving inference accuracy.
3.2 Key-Value Caching in Transformer Models
Key-Value (KV) caching is a well-established optimization in autoregressive Transformers (Vaswani, 2017; Floridi & Chiriatti, 2020), allowing previously computed key ($\mathbf{K}$) and value ($\mathbf{V}$) representations to be stored and reused in subsequent timesteps. Given input tokens $\mathbf{X}$, self-attention first projects them into queries, keys, and values:
$\mathbf{Q}=\mathbf{X}W_{Q},\quad\mathbf{K}=\mathbf{X}W_{K},\quad\mathbf{V}=\mathbf{X}W_{V}.$ (1)
The attention output is computed as:
$\mathrm{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{D}}\right)\mathbf{V}.$ (2)
In standard inference, new keys and values are appended to the cache:
$\mathbf{K}_{t}=\mathrm{Concat}(\mathbf{K}_{t-1},\mathbf{K}_{\text{new}}),\quad\mathbf{V}_{t}=\mathrm{Concat}(\mathbf{V}_{t-1},\mathbf{V}_{\text{new}}).$ (3)
While this improves efficiency by avoiding redundant recomputation, it does not reduce memory or computational costs, as all tokens are still processed.
VLA-Cache extends this approach by introducing dynamic token reuse. Instead of appending all new tokens, it identifies static tokens across frames and skips their recomputation:
$\mathbf{K}_{t}(i)=\begin{cases}\mathbf{K}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}},\\ W_{K}\mathbf{H}_{t}(i),&\text{otherwise}.\end{cases}$ (4)
$\mathbf{V}_{t}(i)=\begin{cases}\mathbf{V}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}},\\ W_{V}\mathbf{H}_{t}(i),&\text{otherwise}.\end{cases}$ (5)
where $\mathcal{P}_{\text{reuse}}$ represents the subset of tokens that remain visually and semantically unchanged between frames. By selectively updating only task-relevant tokens, VLA-Cache significantly reduces redundant computation while preserving high-fidelity action predictions.
This caching mechanism integrates seamlessly into the Transformer decoding process, requiring no architectural modifications or additional training. The following sections detail the implementation of VLA-Cache and its effectiveness in improving inference efficiency for VLA models.
Figure 2: In this paper, we propose VLA-Cache, which includes two procedures. (a) Dynamic Token Selection: First, a static token set is identified, followed by filtering out task-relevant tokens to ensure critical information undergoes full computation. Tokens that remain in the set and overlap with those from the previous step will be retrieved from cache. (b) Adaptive Token Caching: The fraction of reused tokens is dynamically adjusted at each layer based on its attention distribution.
4 Methodology
In robotic action prediction, consecutive image frames often show only minor changes in the manipulator or target objects, while the background remains mostly unchanged. This motivates us to reuse some of unchanged tokens in the last step to accelerate the inference at current step. However, the choice of reusing tokens should be made very carefully, as even small changes to the manipulator or target objects can be pivotal for accurate action prediction. Treating such tokens as static could severely degrade performance. Consequently, the core challenge in accelerating VLA inference via token caching lies in accurately distinguishing genuinely static regions from those that are dynamic or closely tied to the task. By automatically detecting static tokens across consecutive frames and discarding those related to the manipulator or the immediate task, one can realize substantial computational savings while still preserving high-quality action predictions.
4.1 Dynamic Token Selection
Static Token Selection. To enhance inference efficiency, we identify visually static tokens across consecutive frames and reuse their associated Key-Value (KV) representations. This process involves three steps: partitioning the input image into patches, computing patch-wise similarity, and selecting the most static patches for reuse.
Given an image $I$ of size $H\times W$, we divide it into $N\times N$ non-overlapping patches of size $p\times p$, yielding $N^{2}$ patches $\mathcal{P}=\{\mathbf{P}_{i,j}\}$. Each patch $\mathbf{P}_{i,j}$ is directly represented by its raw pixel values, without additional feature encoding.
For two consecutive frames $I_{t}$ and $I_{t-1}$, we calculate the cosine similarity between their corresponding patches:
$\text{Sim}\big{(}\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j}\big{)}=\frac{\mathbf{P}_{t}^{i,j}\cdot\mathbf{P}_{t-1}^{i,j}}{\|\mathbf{P}_{t}^{i,j}\|_{2}\,\|\mathbf{P}_{t-1}^{i,j}\|_{2}}.$ (6)
A patch is considered static if its similarity score exceeds a predefined threshold $\tau$. To further refine the selection, we apply a Top-$k$ filter:
$\mathcal{P}_{\mathrm{top\text{-}k}}=\mathrm{Top}\text{-}k\Big{(}\{\mathbf{P}_{t}^{i,j}\mid\mathrm{Sim}(\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j})\geq\tau\}\Big{)}.$ (7)
Thus, patches with minimal visual variation are identified for reuse, while those displaying notable differences are recomputed. This approach accurately selects truly static tokens across consecutive frames, significantly reducing redundant computations and accelerating inference without compromising overall performance.
Evicting Task-Relevant Tokens. While many tokens remain visually static across consecutive frames, certain tokens that are highly relevant to the task (e.g., the gripper or the target object) cannot be directly reused. In our experiments, we observed that as the robotic arm approaches the target object, small but critical changes in the target region often occur. These changes are primarily caused by factors such as lighting variations or camera noise. Recomputing tokens around the target region before the gripper reaches it significantly improves task accuracy. To address this issue, we leverage the attention mechanism of the VLA model to identify task-relevant tokens in the next frame that require recalculation.
Specifically, we calculate the text-to-vision attention scores from the cross-attention module of the VLA model to determine the task relevance of each vision token. For each layer $l$, the attention weights are represented as $\mathbf{A}^{l}\in\mathbb{R}^{N_{\text{heads}}\times N_{\text{tokens}}\times N_{\text{tokens}}}$. From these weights, we extract the attention mapping from text tokens to vision tokens:
$\mathbf{A}^{l}_{\text{vis-text}}=\mathbf{A}^{l}[:,\,v_{\text{start}}:v_{\text{end}},\,t_{\text{start}}:t_{\text{end}}],$ (8)
where $v_{\text{start}},v_{\text{end}}$ and $t_{\text{start}},t_{\text{end}}$ are the indices of the vision and text tokens, respectively.
To aggregate the attention scores across multiple heads, we compute the mean attention for each vision token as $\mathbf{A}^{l}_{\text{avg}}=\mathrm{Mean}_{\text{heads}}\!\bigl{(}\mathbf{A}^{l}_{\text{vis-text}}\bigr{)}$. For task relevance across multiple layers $\mathcal{L}$, the final task relevance scores are obtained by averaging the scores across the selected layers as $\mathbf{S}_{\text{task-relevance}}=\mathrm{Mean}_{l\in\mathcal{L}}\!\bigl{(}\mathbf{A}^{l}_{\text{avg}}\bigr{)}$.
Using these scores, we rank the vision tokens based on their task relevance and apply a threshold $\tau_{\text{task}}$ to select the most task-relevant tokens:
$\mathcal{P}_{\text{task-relevant}}=\{\mathbf{P}_{t}^{i,j}\mid\mathbf{S}_{\text{task-relevance}}[i,j]\geq\tau_{\text{task}}\}.$ (9)
Finally, we combine the set of static tokens $\mathcal{P}_{\mathrm{static}}$ selected in the first step with the task-relevant tokens. Tokens that are both static and highly task-relevant are removed from the reusable token set to ensure they are recomputed in the current step:
$\mathcal{P}_{\mathrm{final}}=\mathcal{P}_{\mathrm{static}}\;\setminus\;\mathcal{P}_{\text{task-relevant}}.$ (10)
By filtering out high-attention tokens, this approach ensures that the VLA model accurately updates the critical objects in the scene, thereby preventing failures caused by reusing outdated static tokens. This method effectively balances computational efficiency with task precision, enabling robust performance in robotic control tasks.
4.2 Layer Adaptive Token Reusing
While static token selection and task-relevance filtering eliminate a large portion of redundant computation, we observe that attention distributions within the VLA decoder vary significantly across different layers. This finding is consistent with observations reported by prior work FastV (Chen et al., 2025), indicating that both VLA and VLM decoders exhibit similar patterns of attention flow: early layers display dispersed attention, followed by fluctuations in intermediate layers, and eventually a partial rebound near the final layers.
To account for these differences, we propose a layer-adaptive strategy that adjusts the fraction of reused tokens based on each layer’s attention concentration. Specifically, we quantify the attention distribution at layer $l$ via an entropy measure, following the same mean-attention computation described in equation  8. Let $H^{l}$ denote the resulting entropy. We then define an entropy ratio $R^{l}$, which captures how much more concentrated the attention is in layer $l$ compared to layer $l-1$:
$R^{l}\;=\;\frac{H^{l-1}-H^{l}}{H^{l-1}}.$ (11)
A positive $R^{l}$ indicates that the attention distribution at layer $l$ is more focused than that of layer $l-1$.
We accumulate these ratios across layers to obtain a cumulative score $R_{\text{cum}}^{l}=\sum_{j=1}^{l}R^{j}$, which in turn determines the proportion $\alpha^{l}$ of static tokens (from $\mathcal{P}_{\mathrm{final}}$) that are reused at layer $l$. Formally,
[…]
where $k$ is a hyperparameter that governs the impact of attention concentration. Layers with larger cumulative entropy reduction are allowed to reuse a higher fraction of tokens, reflecting the insight that as attention becomes more focused, fewer tokens are likely to require recomputation.
In practice, this layer-adaptive mechanism dynamically adjusts token reuse based on the evolving attention patterns in the VLA decoder, effectively balancing computational efficiency with task accuracy. By selectively retaining only the most relevant tokens at each layer, our method significantly reduces redundant computations while maintaining reliable action prediction.
Algorithm 1 Adaptive Token Caching
1:  Input: Frames $\{I_{t-1},I_{t}\}$, previous KV cache $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$, thresholds $\tau,\tau_{\text{task}}$, hyperparameter $k$
2:  Output: Updated KV cache $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
3:  Static Token Selection:
4:  Patchify $I_{t-1},I_{t}$ and compute $\text{Sim}(\mathbf{P}_{t},\mathbf{P}_{t-1})$
5:  Select $\mathcal{P}_{\text{static}}$ where similarity $\geq\tau$
6:  Apply top-$k$ filtering to refine static token selection
7:  Evict Task-Relevant Tokens:
8:  Compute text-to-vision attention scores $\mathbf{S}_{\text{task-relevance}}$
9:  Select $\mathcal{P}_{\text{task-relevant}}$ where attention $\geq\tau_{\text{task}}$
10:  Compute reusable tokens: $\mathcal{P}_{\mathrm{final}}=\mathcal{P}_{\text{static}}\setminus\mathcal{P}_{\text{task-relevant}}$
11:  Layer-Adaptive Token Reuse:
12:  Compute entropy reduction $R^{l}$, $R_{\text{cum}}^{l}$; Determine reuse ratio $\alpha^{l}$; Compute reusable subset $\mathcal{P}_{\mathrm{reuse}}\subseteq\mathcal{P}_{\mathrm{final}}$
13:  for each layer $l$ and token $i$ do
14:     if $i\in\mathcal{P}_{\mathrm{reuse}}$ then
15:        Reuse cached values:    $\mathbf{K}_{t}^{l}(i)=\mathbf{K}_{t-1}^{l}(I),\mathbf{V}_{t}^{l}(i)=\mathbf{V}_{t-1}^{l}(i)$
16:     else
17:        Recompute:    $\mathbf{K}_{t}^{l}(i)=W_{K}^{l}\mathbf{H}_{t}^{l}(i),\mathbf{V}_{t}^{l}(i)=W_{V}^{l}\mathbf{H}_{t}^{l}(i)$
18:     end if
19:     Compute self-attention:
20:     $\mathrm{Attn}(\mathbf{Q}_{t}^{l},\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l})=\mathrm{Softmax}\!\Bigl{(}\frac{\mathbf{Q}_{t}^{l}(\mathbf{K}_{t}^{l})^{\top}}{\sqrt{D}}\Bigr{)}\mathbf{V}_{t}^{l}$
21:  end for
22:  return Updated KV cache $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
Inference with Cached Representation. During inference, VLA-Cache selectively caches and reuses previously computed representations to reduce redundant computations while maintaining accurate robotic action prediction. Instead of recomputing all tokens at each timestep, the model distinguishes between static and dynamic tokens. Static tokens, which exhibit minimal visual change across consecutive frames, retain their cached key-value (KV) representations from the previous timestep, bypassing unnecessary computation. Dynamic tokens, which undergo significant changes or are task-relevant, are recomputed to ensure precise motion execution. This selective token reuse is performed at each layer, guided by entropy-based attention analysis to dynamically adjust reuse ratios. The full inference procedure, including static token selection, task-relevance filtering, and layer-adaptive token reuse, is outlined in Algorithm 1, with further details provided in Appendix A.2.
4.3 Theoretical Analysis of Computational Complexity
Computational Cost Reduction. In standard VLA inference, each Transformer layer processes $L$ tokens, with a total FLOP cost per layer:
$\text{FLOPs}\approx 4LD^{2}+2L^{2}D+2LDM.$ (13)
Our method reduces the effective token count per layer to $L_{r}=\alpha\times\mathcal{P}_{\mathrm{final}}$, leading to theoretical savings:
$\Delta\text{FLOPs}_{\mathrm{layer}}\approx 4L_{r}D^{2}+2L_{r}^{2}D+2L_{r}DM.$ (14)
These savings scale across multiple layers, significantly lowering computation while maintaining accuracy.
Overhead of Token Selection. The cost of static token identification is approximately $\mathcal{O}(H^{2})$ due to patch similarity checks, while task-relevance filtering introduces a cross-modal attention aggregation cost of $\mathcal{O}(L_{t}L_{v}D)$. The entropy-based layer-adaptive strategy incurs an additional $\mathcal{O}(L^{2}D)$ complexity, which remains significantly lower than the baseline per-layer cost.
Total Complexity Reduction. Bringing all components together, the theoretical overall FLOP reduction per layer is:
$\Delta\text{FLOPs}_{\text{total}}\approx$ $\Bigl{(}4L_{r}D^{2}+2L_{r}^{2}D+2L_{r}DM\Bigr{)}$ (15)
$-\Bigl{(}H^{2}+L_{t}L_{v}D+L^{2}D\Bigr{)}.$
Further Derivations and Complexity Details. See Appendix A.1 for detailed derivations, including static token selection costs, attention filtering complexity, and layer-adaptive entropy calculations.
5 Experiment
5.1 Experimental Settings
To validate the generalization of VLA-Cache, we implement our method in both simulation and real-world settings. In simulation, we evaluate VLA-Cache on two open-source VLA models: OpenVLA (Kim et al., 2024) and CogAct (Li et al., 2024a), using the LIBERO benchmark (Liu et al., 2024a) and SIMPLER environment (Li et al., 2024b), respectively. For fair comparison, we adopt two token-pruning acceleration methods, SparseVLM (Zhang et al., 2024) and FastV (Chen et al., 2025), as baselines in the LIBERO environment. OpenVLA processes 256 visual tokens per image, and all comparative experiments use an identical token pruning and reuse count, setting Top-K to 100 tokens.
For VLA-Cache, the default parameters are set as follows: similarity score threshold $\tau=0.996$, top-$k=100$, and task relevance threshold $\tau_{task}=0.5$. Additionally, we use OpenVLA’s default caching mechanism and other default parameters for inference acceleration. In the SIMPLER environment, we apply VLA-Cache with the same parameter settings as OpenVLA while using CogAct’s default hyperparameters.
For real-world evaluation, we fine-tune OpenVLA for deployment on a Kinova Jaco2 6-DoF manipulator. The model is trained for 50,000 steps with default parameters. During testing, we modify only the similarity score threshold to $\tau=0.85$, keeping all other parameters unchanged. All experiments are conducted on an NVIDIA RTX 4090 GPU.
5.1.1 LIBERO Benchmark
The LIBERO Benchmark (Liu et al., 2024a) is designed to evaluate lifelong robotic manipulation across four diverse task suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each suite introduces distinct variations in spatial layouts, object selection, task objectives, or long-horizon planning, challenging models to generalize across different manipulation scenarios.
Following the experimental setup of OpenVLA, we conduct evaluations on all four LIBERO suites, each containing ten subtasks. To ensure consistency, we use the same machine and official fine-tuned weights. For each subtask, multiple evaluation episodes are run, measuring key performance metrics such as success rate and inference latency. We evaluate VLA-Cache’s acceleration capabilities within the OpenVLA framework and conduct an ablation study on the top-$k$ parameter in the LIBERO environment.
Figure 3: Simulation Tasks on LIBERO Benchmark and the SIMPLER Environment.
Figure 4: Robot Manipulation Tasks on Jaco2 Real-world Platform.
Figure 5: Heatmap Visualization of Robot Actions in a Simulated Environment with Prompt Descriptions and Outputs
[…]
LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
Baseline (OpenVLA) 84.4% 86.6% 75.6% 53.2% 75.0% 1.864 51.91
SparseVLM 79.8% 67.0% 72.6% 39.4% 64.7% 1.407 83.39
FastV 83.4% 84.0% 74.2% 51.6% 73.3% 1.864 53.28
VLA-Cache 83.8% 85.8% 76.4% 52.8% 74.7% 1.355 31.83
Table 1: Comparison of different VLA acceleration methods on the LIBERO benchmark.
[…]
PickCan MoveNear Drawer DrawerApple Average
Visual Matching Baseline (CogACT) 91.3% 85.0% 71.8% 50.9% 74.8% 1.847 54.29
VLA-Cache 92.0% 83.3% 70.5% 51.6% 74.4% 1.496 39.63
Variant Aggregation Baseline (CogACT) 89.6% 80.8% 28.3% 46.6% 61.3% 1.807 53.54
VLA-Cache 91.7% 79.3% 32.5% 45.8% 62.3% 1.493 39.11
Table 2: Comparison of VLA Cache in CogACT model within the SIMPLER environment.
5.1.2 SIMPLER Environments
The SIMPLER simulation environment (Li et al., 2024b) is designed to closely mirror real-world robotic setups, facilitating more realistic evaluations.
SIMPLER supports two evaluation settings: Visual Matching (aiming to minimize visual discrepancies between simulation and reality) and Variant Aggregation (building on Visual Matching by introducing variations in lighting, background, table textures, etc.). We adopt the same SIMPLER configuration used by CogAct and evaluate two settings, Visual Matching and Variant Aggregation, on a Google robot arm. Each setting contains four tasks: 1) Pick coke can(PickCan), 2) Move near(MoveNear), 3) Open/close drawer(Drawer), 4) Open top drawer and place apple(DrawerApple).
To validate the general applicability of our method, we use CogAct as the baseline model within the SIMPLER environment. CogAct extends OpenVLA by incorporating a Diffusion Policy head to predict actions in continuous space, a common strategy among recent VLA models. Similar to OpenVLA, CogAct provides publicly available weights and code for reproducible simulation-based evaluations. Through these experiments, we try to verify its potential as a general VLA acceleration technique.
5.1.3 Real-world Robot Setting
We further evaluate our method on a Kinova Jaco2 6-DoF manipulator, equipped with a Sony AX53 camera positioned in front of the arm to capture the manipulation scene.
Tasks. We design four real-world robotic manipulation tasks:1) Pick up the orange pot (PickPot),2) Place the blue cube in the box (PlaceCube), 3) Put the sausage in the blue pan (PutSausage),4) Wipe the table (WipeTable). For the wiping task, we place objects of various shapes and colors on the table to challenge the model’s robustness to unseen objects.
Data Collection. Using an Xbox Gamepad, we teleoperate the Jaco2 arm and collect demonstrations at a frequency of 10 Hz, recording RGB images and robot states throughout each trajectory. We gather 150–200 demonstrations per task to form the training dataset.
Baseline. We fine-tune OpenVLA via LoRA (Low-Rank Adaptation (Hu et al., 2021)) on each task and then deploy it on the Jaco2 to assess real-world performance. The fine-tuning process follows the same hyperparameter setup as in LIBERO.
5.2 Results on Simulation Environment
[…]
PickPot PlaceCube PutSausage WipeTable Average
Baseline (OpenVLA) 95.0% 83.3% 80.0% 70.0% 82.1% 1.814 64.16
VLA-Cache 90.0% 90.0% 85.0% 73.0% 84.5% 1.303 51.85
Table 3: Comparison of success rate on real robot tasks.
[…]
Static Token 74.2% 1.304 31.03
Task Relevant 82.6% 1.304 31.03
Layer Adaptive (Ours) 83.8% 1.382 32.22
Table 4: Comparison of different token selection methods.
  #Tokens Methods SR % $\uparrow$ FLOPs $\downarrow$ Time(ms) $\downarrow$
0 Baseline 84.4 1.888 52.37
50 SparseVLM 79.8 1.358 88.08
FastV 84.6 1.888 53.10
Ours 85.4 1.611 33.43
100 SparseVLM 74.6 1.097 61.01
FastV 83.4 1.888 45.72
Ours 83.8 1.295 31.29
200 SparseVLM 44.4 0.735 57.42
FastV 72.8 1.888 45.19
Ours 68.3 0.823 30.29
Table 5: LIBERO-Spatial benchmark comparing different methods based on pruning or reuse tokens. Best values in each column are highlighted in bold.
Main Results on LIBERO.
Table 1 summarizes the success rate, FLOPs, and CUDA inference time across the four LIBERO task suites. Compared to the baseline, VLA-Cache achieves a 27.31% FLOPs reduction and a 1.63$\times$ speedup, with only a minimal 0.3% drop in the overall success rate, confirming its efficiency without compromising accuracy. While VLA-Cache slightly underperforms on spatial, object, and long-horizon tasks, it notably surpasses the baseline on the goal suite, highlighting its robust adaptability.
Notably, both FastV and SparseVLM fail to improve inference time, often exceeding the baseline. FastV merely masks tokens in attention computation without reducing GPU workload, introducing additional masking overhead, while SparseVLM incurs extra cost from token pruning, merging, and recycling. These methods were originally designed for long-sequence VLM tasks, where efficiency gains are more pronounced, but struggle in VLA settings, where action sequences are typically short. In contrast, VLA-Cache leverages inter-frame token reuse with negligible overhead, making it effective for robotic action generation.
Ablation on Reusing/Pruning Rate. Table 5 presents a comparative analysis of different methods through an ablation study on the top-$k$ parameter, while also varying the number of pruned or reused tokens. For all methods, increasing the number of reused/pruned tokens leads to a decline in success rates, highlighting the importance of retaining sufficient information for accurate action predictions. Our approach maintains a relatively stable success rate at a reuse rate of 40%, but a more noticeable drop occurs when reusing 200 tokens. In contrast, FastV removes a significant portion of tokens entirely, and its performance drop primarily results from the loss of critical visual details rather than potential errors in token reuse. Additionally, SparseVLM and FastV incur longer inference times due to extra attention operations during GPU execution, whereas VLA-Cache directly updates KV cache entries, making inference more efficient.
Token Selection Strategies As presented in Table 4, we evaluate different token selection strategies for VLA-Cache. The direct reuse of static tokens results in a significant decline in success rate. In contrast, filtering task-relevant tokens maintains a success rate of 82.6%. Notably, our layer-adaptive approach further enhances the success rate to 83.8%, albeit with a marginal increase in computational overhead from 1.304 FLOPs to 1.382 FLOPs. This overhead arises because fewer tokens are reused in the early layers, which ensures the retention of critical updates in deeper layers, thereby improving overall performance.
Main Results on SIMPLER
Table 2 indicates that the VLA-Cache exhibits comparable success rates than the baseline CogACT in the SIMPLER environment while substantially reducing computational overhead. Specifically, VLA-Cache achieves an average success rate of 74.4% in the standard Visual Matching setting compared to 74.8% for CogACT and 62.3% in the Variant Aggregation setting versus 61.3% for CogACT.
The efficiency gains are evident in the FLOPs and inference time measurements. VLA-Cache achieves roughly 20% fewer FLOPs than the baseline, coupled with a 1.37× reduction in inference latency. Notably, these results highlight the portability of VLA-Cache across different action heads, establishing it as a general acceleration strategy for VLA.
5.3 Results on Real Robot
Table 3 illustrates the performance of VLA-Cache in real-world robotic tasks. Among the four tasks, PickPot shows a slightly lower success rate than the baseline, whereas VLA-Cache exceeds the baseline on the other three tasks. The method also achieves considerable reductions in FLOPs and inference time. Overall, VLA-Cache improves the average success rate by 2.4%, an outcome that, although unexpected, is potentially explained by a more robust baseline model, which we trained extensively using additional data and more training steps than the LIBERO fine-tuning. With improved robustness as a foundation, VLA-Cache’s ability to prune or reuse redundant tokens may further enhance the model’s resilience, thus yielding higher success rates.
6 Conclusion
In this paper, we introduce VLA-Cache, a training-free method for VLA that selectively reuses static tokens while filtering out task-relevant ones, reducing redundant computation without sacrificing accuracy. Additionally, our layer-adaptive token reuse strategy improves model success rates by adjusting token reuse based on attention concentration. Extensive experiments on two VLA models, OpenVLA and CogAct, across two simulation environments, LIBERO and SIMPLER, demonstrate that VLA-Cache achieves a 1.7$\times$ speedup while maintaining performance.
Appendix A Appendix
A.1 Complexity Analysis Details
Static Token Selection. We compute patch-wise similarity for visual tokens:
$\text{FLOPs}_{\text{static-sim}}=N_{\mathrm{patch}}^{2}D_{\mathrm{patch}}\approx H^{2}.$ (16)
Since $N_{\mathrm{patch}}=H/p$, this cost remains small relative to Transformer computations.
Task-Relevance Filtering. The attention-based filtering step computes cross-modal importance scores:
$\text{FLOPs}_{\text{task-filter}}\approx L_{t}L_{v}D.$ (17)
A sorting operation of $\mathcal{O}(L\log L)$ follows for threshold selection.
Layer-Adaptive Entropy Computation. The entropy-based reuse strategy involves:
$\text{FLOPs}_{\text{entropy}}\approx L^{2}D.$ (18)
The per-layer reuse ratio is then computed as:
$\alpha^{l}=\min\Bigl{(}k\sum_{j=1}^{l}R^{j},1\Bigr{)}.$ (19)
These overheads remain modest compared to the full forward pass cost, enabling efficient token reuse.
Final FLOP Reduction. The total FLOP savings across all layers follow:
$\Delta\text{FLOPs}_{\text{total}}=\sum_{l=1}^{\Omega}\Delta\text{FLOPs}_{\text{layer}}.$ (20)
This confirms that dynamic token reuse significantly reduces computation without sacrificing model performance.
A.2 VLA Cache Inference
Having identified both static tokens and task-relevant tokens as described in Section 4.1, we now detail how these tokens are reused in the VLA decoder via KV caching. Below, we first present a high-level overview of the procedure, then introduce several formulas to clarify the partial update of key-value (KV) representations under our selective token reuse strategy.
Overall Procedure.
At time step $t-1$, while predicting the action for frame $I_{t-1}$, the model produces the attention weights (used to evaluate task relevance) and the KV cache, $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$ for each layer $l$. When a new frame $I_{t}$ arrives, we compare it with $I_{t-1}$ (6) to obtain a set of static tokens sorted by similarity scores, and further filter out any high task-relevance tokens (9). Simultaneously, we determine a layer-specific reuse ratio $\alpha^{l}$  (12) based on the entropy of attention distributions at time $t-1$. This yields:
$\mathcal{P}_{\mathrm{final}}\;\text{and}\;\{\alpha^{l}\}_{l=1}^{\Omega},$ (21)
where $\mathcal{P}_{\mathrm{final}}$ denotes the final subset of tokens eligible for reuse, and $\Omega$ is the total number of decoder layers.
Selective Token Reuse.
During decoding at time $t$, each decoder layer $l$ reuses only the top-$k$ tokens (determined by $\alpha^{l}$ and similarity ranking) from $\mathcal{P}_{\mathrm{final}}$. Let $\mathcal{P}_{\mathrm{reuse}}\subseteq\mathcal{P}_{\mathrm{final}}$ be this chosen set of reusable tokens, with positions (IDs) $\{i\mid i\in\mathcal{P}_{\mathrm{reuse}}\}$. For tokens not in $\mathcal{P}_{\mathrm{reuse}}$, we recompute their key-value (KV) representations from scratch. Concretely, let $\mathbf{H}_{t}^{l}\in\mathbb{R}^{L\times D}$ be the hidden states at layer $l$ for the new frame $I_{t}$. The per-layer self-attention requires three components: queries $\mathbf{Q}_{t}^{l}$, keys $\mathbf{K}_{t}^{l}$, and values $\mathbf{V}_{t}^{l}$. We define:
$\mathbf{Q}_{t}^{l}(i)\;=\;W_{Q}^{l}\,\mathbf{H}_{t}^{l}(i),$ (22)
for all token indices $i\in\{1,\ldots,L\}$, where $W_{Q}^{l}\in\mathbb{R}^{D\times D}$ is the learnable projection of queries.
For $\mathbf{K}_{t}^{l}$ and $\mathbf{V}_{t}^{l}$, we partially reuse cached values from time $t-1$:
$\mathbf{K}_{t}^{l}(i)\;=\;\begin{cases}\mathbf{K}_{t-1}^{l}(i),&\text{if }i\in\mathcal{P}_{\mathrm{reuse}},\\ W_{K}^{l}\,\mathbf{H}_{t}^{l}(i),&\text{otherwise},\end{cases}$ (23)
$\mathbf{V}_{t}^{l}(i)\;=\;\begin{cases}\mathbf{V}_{t-1}^{l}(i),&\text{if }i\in\mathcal{P}_{\mathrm{reuse}},\\ W_{V}^{l}\,\mathbf{H}_{t}^{l}(i),&\text{otherwise},\end{cases}$ (24)
where $W_{K}^{l},W_{V}^{l}\in\mathbb{R}^{D\times D}$ are the key and value projection matrices, respectively, and $\mathbf{K}_{t-1}^{l}(i),\mathbf{V}_{t-1}^{l}(i)$ are the cached key-value vectors from the previous time step. This partial-update scheme allows us to skip computation for those tokens deemed both static and non-task-critical.
Finally, the self-attention output at layer $l$ is computed as:
$\mathrm{Attn}(\mathbf{Q}_{t}^{l},\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l})\;=\;\mathrm{Softmax}\!\Bigl{(}\tfrac{\mathbf{Q}_{t}^{l}(\mathbf{K}_{t}^{l})^{\top}}{\sqrt{D}}\Bigr{)}\;\mathbf{V}_{t}^{l}.$ (25)
Only the newly computed tokens update $\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}$, while reused tokens retain their entries from $\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}$.
Implementation Details.
Listing A.2 outlines our partial token reuse within the VLA decoder:
  1. 1.
    Position IDs and Causal Masks. We maintain a cache_position array reflecting the positions of tokens that require computation. Tokens not recomputed do not receive new position IDs, enabling a pruned causal mask aligned with the updated token set.
  2. 2.
    Computing Rotary Embeddings. Once the reusable tokens are omitted from re-encoding, we apply a rotary embedding function to the reduced set of token embeddings to generate position-dependent features for the next layer.
  3. 3.
    KV Cache Updates. Only newly computed tokens update their corresponding key-value entries in $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$. Reused tokens keep their entries from $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$. This partial-update mechanism leverages the permutation-invariance property of Transformers, ensuring consistent attention outputs even if certain tokens are omitted from recomputation.
Since Transformers naturally support skipping previously processed tokens in a self-attention cache, our layer-adaptive token reuse is readily integrated with standard caching strategies. Notably, the most significant reduction in computation occurs when generating the first action token at time $t$. After that, the autoregressive procedure continues as usual, incurring no additional overhead for subsequent tokens.
Efficiency and Effect on Performance.
By leveraging selective token reuse, VLA-Cache decreases the total number of tokens that must be re-encoded and re-attended at each time step $t$. This leads to a substantial reduction in decoding time, especially given the short sequences in robotic action prediction. Meanwhile, the complexity of predicting subsequent tokens in the same time step follows the usual KV-caching paradigm, adding minimal overhead. In practice, this method preserves task accuracy by forcing recomputation of highly dynamic or task-critical tokens while reusing the remainder. We present a detailed evaluation in Section 5, demonstrating that VLA-Cache achieves notable speedups with negligible performance degradation.
A.3 Finetuning Data for Real Robot Experiments
Robot Setup.
The setup of the Franka Robot is shown in Figure  6. In this example, a Kinova Jaco robot arm with 6 degrees of freedom is rigidly fixed to the frame. We use a Sony AX53 camera, which is placed opposite the robot arm. The camera is facing the operating table and transmits the video in real time.
Figure 6: Kinova Jaco Robot Setup
Data Collection.
Our data collection work is based on the CLVR_Jaco_Play dataset. We employed PyBullet as an inverse kinematics (IK) controller. The system receives incremental Cartesian displacement inputs ($\Delta x,\Delta y,\Delta z$) from an Xbox controller, which are processed to generate joint velocity commands. These commands are transmitted to the robotic arm's velocity controller at a 10Hz control frequency for real-time execution. We record four types of observations: Third-person camera observations (front_cam_ob), End-effector Cartesian pose (ee_cartesian_pos_ob), End-effector Cartesian velocity (ee_cartesian_vel_ob), Jaco arm joint positions (joint_pos_ob).
Data Preprocessing.
The preprocessing procedure was conducted as follows: Initially, the recorded frames underwent center cropping, reducing the resolution from 1280×720 to 912×720, followed by resizing to 224×224. Subsequently, episodes were manually selected based on a visual inspection of the video sequences generated from the recorded frames. To minimize potential biases associated with excessively long episodes, those exceeding 250 steps were excluded from the dataset. Furthermore, steps in which all recorded action values were zero were removed to ensure data relevance and integrity.
A.4 Simulated Evaluation Detail.
LIBERO Task Definitions.
Similarly, we also utilize all task suites provided in LIBERO for our evaluations. The Robosuite-based robot setup includes the following tasks: 1) “place bowl on plate with spatial variation” (e.g., drawer positions), 2) “pick object” (e.g., ketchup, bowl, apple), 3) “(open / close) target drawer; action object” (e.g., “open top drawer; place apple into drawer”), and 4) “achieve goal using shared objects” (e.g., rearranging spatial relationships or altering object states).
SIMPLER Task Definitions.
We utilize all task variants provided in SIMPLER for our evaluations, which include the Google robot setup with the following tasks: 1) “pick Coke can”, 2) “move obj1 near obj2”, 3) “(open / close) (top / middle / bottom) drawer”, and 4) “open top drawer; place apple into top drawer”. Evaluations for the Google robot setup are provided for both Visual Matching (VM) and Variant Aggregations (VA).
Implementation Details.
Simulated evaluations for CogACT and SIMPLER are conducted on a single NVIDIA RTX 4090 GPU in BF16 precision. During inference, we use DDIM sampling with 10 steps and a classifier-free guidance (CFG) coefficient of 1.5. Similarly, for OpenVLA and LIBERO, inference is performed on a single NVIDIA RTX 4090 GPU in BF16 precision.
A.5 Additional Simulation Results
Result of Subtask on LIBERO Spatial Task Suit.
[…]
task1 task2 task3 task4 task5 task6 task7 task8 task9 task10 Avg
Baseline (OpenVLA) 90 90 84 96 70 90 96 76 82 70 84.4 1.888 52.37
SparseVLM 88 60 90 90 60 82 90 92 72 74 79.8 1.367 88.08
FastV 92 90 90 94 58 92 90 80 78 70 83.4 1.888 54.00
VLA-Cache 90 90 88 94 66 84 94 84 76 72 83.8 1.382 32.22
Table 6: Comparison of success rates across different tasks in the LIBERO-Spatial benchmark.
Table 6 presents detailed results on each subtask in the LIBERO-Spatial suite. We observe that VLA-Cache, along with methods like SparseVLM and FastV, occasionally surpasses the baseline’s success rate on individual subtasks. This suggests that certain redundant tokens may distract the baseline model, and pruning or reusing tokens can in fact enhance its robustness.
Figure 7: VLA-Cache test results and attention heat map in a simulated environment
Visualization Results.
Figure 7 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache in LIBERO.
A.6 Real-World Evaluation Detail.
Task Definitions.
  1. 1.
    Pick Up Orange Pot (single-instruction): The robot's goal is to grasp the orange pot and lift it completely off the table. We collected 218 valid demonstrations for the training dataset, with slight random adjustments to the initial positions of the pot and the robotic arm in each episode. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  2. 2.
    Place Blue Cube in Box (single-instruction): The robot's goal is to place the held blue cube into the target container. We collected 212 valid demonstrations for the training dataset, with randomized container positions and robotic arm initial configurations in each episode. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  3. 3.
    Put Sausage in Blue Pan (single-instruction): The robot's goal is to stably place the held sausage into the blue pan. We collected 219 valid demonstrations for the training dataset, with randomized pan coordinates and robotic arm joint angles across trials. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
  4. 4.
    Wipe Table (single-instruction): The robotic arm's goal is to sweep scattered items into a fixed-position dustpan using a broom. For the training dataset, we collected 187 valid demonstrations featuring randomized placements of simulated items (e.g., fries/cheese) on the table and varied initial joint configurations of the robotic arm, while the dustpan location remained fixed. This task explicitly incorporates dynamic environmental variations to rigorously test generalization capabilities. During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
Visualization Results.
Figure 8 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache on the Kinova Jaco Robot Arm.
Figure 8: Comparison between baseline(OpenVLA) and VLA-Cache in real environment

### v2

@@
-VLA-Cache: Towards Efficient Vision-Language-Action Model via Adaptive Token Caching in Robotic Manipulation
+VLA-Cache: Efficient Vision-Language-Action Manipulation via Adaptive Token Caching
 Abstract
-Vision-Language-Action (VLA) model can process instructions and visual perception to directly generate actions as output in an end-to-end fashion due to its strong multi-modal reasoning capabilities.
-While the performance of VLA models is promising, their computational cost can be substantial.
-This raises challenge for applying them on robotics tasks, which requires real-time decision-making to respond quickly to environmental changes.
-Since robotic control involves sequential decision-making, the visual input often exhibits minimal variation between successive steps.
-A natural idea is to reuse the computational results of unchanged visual tokens from the last step.
-Motivated by this idea, we propose VLA-Cache, an efficient vision-language-action model.
-VLA-Cache incorporates a token-selection mechanism that compares the visual input at each step with the input from the previous step, adaptively identifying visual tokens with minimal changes.
-The computational results for these unchanged tokens are then reused in subsequent steps via KV-cache, thereby significantly improving the efficiency of the VLA-Cache model.
-Experimental results on both simulation (e.g., LIBERO benchmark and SIMPLER) and real-world robot valid VLA-Cache can achieve practical acceleration with minimal sacrifice in success rate.
+Vision-Language-Action (VLA) models have demonstrated strong multi-modal reasoning capabilities, enabling direct action generation from visual perception and language instructions in an end-to-end manner.
+However, their substantial computational cost poses a challenge for real-time robotic control, where rapid decision-making is essential.
+This paper introduces VLA-Cache, a training-free inference acceleration method that reduces computational overhead by adaptively caching and reusing static visual tokens across frames.
+Exploiting the temporal continuity in robotic manipulation, VLA-Cache identifies minimally changed tokens between adjacent frames and reuses their cached key-value representations, thereby circumventing redundant computations.
+Additionally, to maintain action precision, VLA-Cache selectively re-computes task-relevant tokens that are environmentally sensitive, ensuring the fidelity of critical visual information.
+To further optimize efficiency, we introduce a layer adaptive token reusing strategy that dynamically adjusts the reuse ratio based on attention concentration across decoder layers, prioritizing critical tokens for recomputation.
+Extensive experiments on two simulation platforms (LIBERO and SIMPLER) and a real-world robotic system demonstrate that VLA-Cache achieves up to 1.7× speedup in CUDA latency and a 15% increase in control frequency, with negligible loss on task success rate.
+The code and videos can be found at our project page: https://vla-cache.github.io.
 1 Introduction
-Learning a robust and generalizable policy for robotic manipulation through policy learning has long been a challenging problem (Kim et al., 2020), with traditional reinforcement learning approaches (Choi et al., 2024; Bica et al., 2021) often suffering from poor robustness and limited generalization.
-Recently, the rapid advancement of foundational Vision-Language Models (VLMs) (Awadalla et al., 2023; Liu et al., 2024b) has demonstrated remarkable capabilities in multimodal understanding and generalization.
-Leveraging large-scale real-world robotic datasets (O’Neill et al., 2024; Fang et al., 2024), pioneering works (Zitkovich et al., 2023; Octo Model Team et al., 2024; Niu et al., 2024; Kim et al., 2024) have introduced Vision-Language-Action (VLA) models, which integrate vision and language modalities to directly generate robotic actions in an end-to-end manner.
+Learning a robust and generalizable policy for robotic manipulation through policy learning has long been a challenging problem [1], with traditional reinforcement learning approaches [2, 3] often suffering from poor robustness and limited generalization.
+Recently, the rapid advancement of foundational Vision-Language Models (VLMs) [4, 5] has demonstrated remarkable capabilities in multimodal understanding and generalization.
+Leveraging large-scale real-world robotic datasets [6, 7], pioneering works [8, 9, 10, 11] have introduced Vision-Language-Action (VLA) models, which integrate vision and language modalities to directly generate robotic actions in an end-to-end manner.
 This emerging paradigm holds great promise for enhancing the adaptability and generalization of robotic control systems, but leaves a large computational demand.
+To mitigate the extensive cost of VLA models, existing works often adopt generic acceleration techniques, such as model lightweighting [12], quantization [13], and early-exit [14].
+While effective to some extent, these methods often require architectural modifications or retraining, and more importantly, they lack task-specific design tailored to the intrinsic characteristics of VLA tasks.
+As a result, they struggle to achieve an optimal balance between inference speed and action accuracy.
+In this paper, we step into the nature of VLA robotic manipulation that is intrinsically different from the VLM models.
+Specifically, VLA tasks involve sequentially processing a stream of temporally adjacent visual observations, where the environment often exhibits high spatial redundancy across time.
+As shown in Figure 1, large portions of the visual scene, especially background regions, remain static and semantically irrelevant to action decisions, yet are processed repeatedly at each time step.
+These static tokens contribute significantly to computational overhead while providing limited utility for downstream control.
+This motivates our proposed token caching mechanism, which explicitly exploits temporal redundancy in visual inputs to reduce redundant computation without compromising decision quality.
 Figure 1: During the inference of the VLA model, static tokens of the input image remain largely consistent across steps.
-This consistency allows for caching the computations of these tokens from the previous step and improving the efficiency of inference.
-To mitigate the extensive cost of VLA models, existing works often adopt generic acceleration techniques, such as model lightweighting (Wen et al., 2024), quantization (Park et al., 2024), and early exit (Yue et al., 2024).
-However, these approaches fail to account for the unique characteristics of VLA models, leading to suboptimal efficiency gains and limited performance improvements.
-In this paper, we argue that robotic manipulation, characterized by sequential action-scene interactions, inherently prioritizes visual inputs related to the robotic arm and the target object, while the background remains largely static and less critical.
-As illustrated in Figure 1, the static tokens of the input image remain largely consistent across steps.
-In this context, the background tokens, which contribute significantly to computational overhead and repeat across sequential steps in the VLA process, are often redundant.
-As a result, we propose VLA-Cache, a novel approach that identifies and reuses static tokens across sequential VLA steps.
-Specifically, static tokens in each step refer to visual tokens that exhibit minor changes compared to the previous step.
-Instead of recomputing these tokens, VLA-Cache directly substitutes them with their cached representations from the preceding step, significantly reducing redundant computations and improving inference efficiency.
-Nevertheless, we observe that model precision is highly sensitive to task-relevant tokens (i.e., those related to the robotic arm and the target object) in the ablation study, where even a slight positional shift can lead to a significant performance drop.
-To address this, we introduce a fine-grained selection scheme that filters out task-relevant tokens from the static token set, ensuring that these critical tokens continue to undergo full computational interactions, thereby preserving model accuracy while maintaining efficiency.
-Additionally, based on the observation that attention distributions within the VLA decoder vary significantly across different layers, we propose a layer-adaptive strategy that adjusts the fraction of reused tokens based on each layer’s attention concentration.
-We evaluate VLA-Cache on a diverse set of robotic manipulation tasks in the LIBERO benchmark, where it demonstrates over 1.7× acceleration compared to the baseline, with only slight drop on the success rate.
-Furthermore, VLA-Cache is deployed on the Kinova Jaco robot arm, showing that it can achieve practical acceleration in real-world scenarios.
+This consistency allows for caching the computations of these tokens from the previous step.
+To address the inefficiency introduced by repeatedly processing static visual information, we present VLA-Cache, a training-free inference acceleration method that exploits temporal continuity in robotic perception.
+Rather than recomputing all vision tokens at every timestep, VLA-Cache identifies tokens that exhibit minimal change between adjacent frames and reuses their cached key-value (KV) representations to bypass redundant computation.
+However, we observe that not all visually static tokens can be safely reused.
+Some tokens, such as those near the gripper or target object, may appear visually unchanged but remain semantically active and crucial for accurate action generation.
+Naively reusing all static tokens results in a significant performance drop as shown in Table  1.
+To mitigate this, VLA-Cache incorporates a lightweight filtering mechanism based on decoder attention scores to exclude task-relevant tokens from reuse, ensuring that semantically critical regions are always recomputed with up-to-date features.
+Moreover, we observe that attention patterns vary across decoder layers, with deeper layers exhibiting more concentrated focus.
+To further optimize reuse, VLA-Cache employs a layer-adaptive caching strategy that dynamically adjusts the reuse ratio per layer based on attention entropy, prioritizing precise updates in sensitive regions.
+These two mechanisms together enable substantial reduction in decoding overhead, especially in large-scale language decoders (e.g., LLaMA [15], Gemma [16]), which typically dominate the compute cost in VLA systems.
+[…]
 2 Related Work
 Vision-Language-Action Models.
-Large-scale vision-language models (VLMs) have significantly advanced multimodal learning by integrating image understanding and language reasoning (Liu et al., 2024b; Bai et al., 2023).
-Extending these capabilities, Vision-Language-Action (VLA) models (Zitkovich et al., 2023; Li et al., 2023) incorporate an action modality, enabling end-to-end visuomotor control.
-These models typically adopt large VLM backbones (Touvron et al., 2023) and fine-tune them on robot data (O’Neill et al., 2024), with approaches varying from discretizing actions as language-like tokens (Kim et al., 2024; Chi et al., 2023) to incorporating specialized diffusion policy heads (Black et al., 2024).
-Despite their effectiveness in tasks like object retrieval and assembly (Huang et al., 2024; Zhao et al., 2023), VLA models demand substantial computation, making real-time deployment challenging, particularly in resource-constrained environments.
+Large-scale vision-language models (VLMs) have significantly advanced multimodal learning by integrating image understanding and language reasoning [5, 21].
+Extending these capabilities, VLA models [8, 22] incorporate an action modality, enabling end-to-end visuomotor control.
+These models typically adopt large VLM backbones [15] and fine-tune them on robot data [6], with approaches varying from discretizing actions as language-like tokens [11, 23] to incorporating specialized diffusion policy heads [24].
+Despite their effectiveness in tasks like object retrieval and assembly [25, 26], VLA models demand substantial computation, making real-time deployment challenging, particularly in resource-constrained environments.
 Acceleration for Vision-Language Models.
-Acceleration techniques for VLMs, such as quantization (Xie et al., 2024) and pruning (Lin et al., 2024a), have shown success in vision-language tasks but often overlook the real-time demands of action generation.
-Token-level methods like FastV (Chen et al., 2025) and SparseVLM (Zhang et al., 2024) optimize inference by pruning or merging redundant tokens, but their effectiveness in VLA tasks, which require short and highly specialized action sequences, remains unclear.
-Some works address efficiency by modifying VLA architectures, such as DeeR-VLA (Yue et al., 2024), which dynamically adjusts inference depth, and QAIL (Park et al., 2024), which integrates quantization-aware training.
-Others, like RoboMamba (Liu et al., 2024c) and TinyVLA (Wen et al., 2024), focus on replacing traditional attention mechanisms or training compact models from scratch, often requiring re-training and additional data collection.
-In contrast, VLA-Cache offers a training-free acceleration approach by selectively caching static tokens and recomputing only dynamic or task-relevant ones, preserving crucial temporal cues while reducing inference latency and improving real-time robotic control efficiency.
-3 Preliminary
-3.1 Vision-Language-Action Models
-Vision-Language-Action (VLA) models extend large-scale VLMs to generate robotic actions based on multimodal inputs, including visual observations and task instructions.
-These models are typically trained on large-scale datasets, such as Open X Embodiment (O’Neill et al., 2024), which contains over a million real-world robot trajectories.
-The training pipeline involves a pre-trained vision encoder (Liu et al., 2021; Radford et al., 2021) extracting image features, which are projected into the input space of a Large Language Model (LLM).
-The LLM, fine-tuned alongside a projector module, encodes these inputs to predict actions.
-During inference, VLA models process sequential frames to iteratively guide the robot, but the full Transformer-based LLM inference at each step results in substantial computational overhead.
-The quadratic complexity of self-attention, combined with large key-value (KV) caches (Ge et al., 2023), leads to high memory consumption and slow inference speeds, posing challenges for real-time deployment.
-While model compression techniques such as quantization (Lin et al., 2024b) and pruning (Park et al., 2024; Yue et al., 2024) have been explored to reduce computational cost, they typically require re-training and may degrade performance, limiting their practicality in real-time applications.
-To mitigate these inefficiencies, our approach leverages token caching mechanisms to accelerate inference in a training-free manner, reducing computational cost while preserving inference accuracy.
-3.2 Key-Value Caching in Transformer Models
-Key-Value (KV) caching is a well-established optimization in autoregressive Transformers (Vaswani, 2017; Floridi & Chiriatti, 2020), allowing previously computed key ($\mathbf{K}$) and value ($\mathbf{V}$) representations to be stored and reused in subsequent timesteps.
-Given input tokens $\mathbf{X}$, self-attention first projects them into queries, keys, and values:
-$\mathbf{Q}=\mathbf{X}W_{Q},\quad\mathbf{K}=\mathbf{X}W_{K},\quad\mathbf{V}=\mathbf{X}W_{V}.$ (1)
-The attention output is computed as:
+Inference acceleration has been extensively explored in vision-language models (VLMs) through quantization [27], pruning [28], and token-level techniques such as FastV [29], SparseVLM [30], ToMe [31], PuMer [32], and MADTP [33].
+These intra-frame strategies reduce redundancy within a single image but disregard the temporal and spatial structure essential for robotic tasks under closed-loop control.
+In the VLA domain, efficiency has been addressed through architectural modifications (e.g., RoboMamba [34], TinyVLA [12]), quantization-aware training (QAIL [13]), and dynamic depth control (DeeR-VLA [14]).
+While effective, these methods require re-training and lack generalizability.
+Recent high-frequency frameworks such as $\pi_{0}$-FAST [35], HiRT [36], and OpenVLA-OFT [20] achieve higher control frequency via action chunking or asynchronous decoding.
+However, they continue to suffer from the language model decoding bottleneck, which dominates inference time.
+VLA-Cache addresses this gap by introducing a cross-frame token reuse strategy that accelerates inference without modifying the model or requiring additional training.
+Furthermore, it complements existing high-frequency VLA architectures by directly accelerating the language decoder bottleneck, offering a lightweight, plug-and-play solution for real-time robotic inference.
+3 Methodology
+In robotic action prediction, most visual tokens remain static across frames except for key regions like the manipulator or target object.
+While this temporal redundancy enables token reuse, reusing all static tokens can harm accuracy when task-relevant regions subtly change.
+To address this, we propose a method that identifies visually static tokens and filters out semantically important ones based on attention scores from the VLA decoder.
+By avoiding redundant computation of unchanged static tokens between adjacent frames, our approach directly alleviates the computation bottleneck of language decoder in VLA models while preserving the accuracy of action prediction.
+3.1 KV Cache for VLA Token Reusing
+Key-Value (KV) caching is a widely adopted technique in large-scale autoregressive models to reduce both computation and memory footprint during decoding.
+Initially proposed in the Transformer architecture [37], KV caching enables the model to reuse previously computed key ($\mathbf{K}$) and value ($\mathbf{V}$) vectors for each token, thereby avoiding redundant computation across decoding steps.
+Concretely, given a sequence of input tokens $\mathbf{X}$, the self-attention mechanism computes:
+$\mathbf{Q}=\mathbf{X}W_{Q},\quad\mathbf{K}=\mathbf{X}W_{K},\quad\mathbf{V}=\mathbf{X}W_{V},$ (1)
 $\mathrm{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}.$ (2)
-In standard inference, new keys and values are appended to the cache:
-$\mathbf{K}_{t}=\mathrm{Concat}(\mathbf{K}_{t-1},\mathbf{K}_{\text{new}}),\quad\mathbf{V}_{t}=\mathrm{Concat}(\mathbf{V}_{t-1},\mathbf{V}_{\text{new}}).$ (3)
-While this improves efficiency by avoiding redundant recomputation, it does not reduce memory or computational costs, as all tokens are still processed.
-VLA-Cache extends this approach by introducing dynamic token reuse. Instead of appending all new tokens, it identifies static tokens across frames and skips their recomputation:
-$\mathbf{K}_{t}(i)=\begin{cases}\mathbf{K}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}},\\ W_{K}\mathbf{H}_{t}(i),&\text{otherwise}.\end{cases}$ (4)
-$\mathbf{V}_{t}(i)=\begin{cases}\mathbf{V}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}},\\ W_{V}\mathbf{H}_{t}(i),&\text{otherwise}.\end{cases}$ (5)
-where $\mathcal{P}_{\text{reuse}}$ represents the subset of tokens that remain visually and semantically unchanged between frames.
-By selectively updating only task-relevant tokens, VLA-Cache significantly reduces redundant computation while preserving high-fidelity action predictions.
-This caching mechanism integrates seamlessly into the Transformer decoding process, requiring no architectural modifications or additional training.
-The following sections detail the implementation of VLA-Cache and its effectiveness in improving inference efficiency for VLA models.
-Figure 2: In this paper, we propose VLA-Cache, which includes two procedures.
-(a) Dynamic Token Selection: First, a static token set is identified, followed by filtering out task-relevant tokens to ensure critical information undergoes full computation.
-Tokens that remain in the set and overlap with those from the previous step will be retrieved from cache.
-(b) Adaptive Token Caching: The fraction of reused tokens is dynamically adjusted at each layer based on its attention distribution.
-4 Methodology
-In robotic action prediction, consecutive image frames often show only minor changes in the manipulator or target objects, while the background remains mostly unchanged.
-This motivates us to reuse some of unchanged tokens in the last step to accelerate the inference at current step.
-However, the choice of reusing tokens should be made very carefully, as even small changes to the manipulator or target objects can be pivotal for accurate action prediction.
-Treating such tokens as static could severely degrade performance.
-Consequently, the core challenge in accelerating VLA inference via token caching lies in accurately distinguishing genuinely static regions from those that are dynamic or closely tied to the task.
-By automatically detecting static tokens across consecutive frames and discarding those related to the manipulator or the immediate task, one can realize substantial computational savings while still preserving high-quality action predictions.
-4.1 Dynamic Token Selection
+During decoding, each new token’s $\mathbf{k}_{\text{new}}$ and $\mathbf{v}_{\text{new}}$ are appended to the $i$-th cache:
+$\mathbf{K}_{i}=\mathrm{Concat}(\mathbf{K}_{i-1},\mathbf{k}_{\text{new}}),\quad\mathbf{V}_{i}=\mathrm{Concat}(\mathbf{V}_{i-1},\mathbf{v}_{\text{new}}).$ (3)
+While KV caching is effective for language decoding within a single query in vision-language models, this technique does not address redundancy in the visual stream, especially in Vision-Language-Action (VLA) models.
+In robotic manipulation, consecutive visual inputs often share large overlapping content, yet VLA models typically discard visual encodings after each step and recompute them from scratch.
+This is both wasteful and suboptimal for real-time control.
+This inefficiency motivates a key question: can we selectively reuse static visual tokens across time with temporal KV Caching in VLA models?
+This idea forms the basis of our proposed VLA-Cache.
+In the following sections, we introduce its core mechanisms: static token selection, task-relevance filtering, and layer-adaptive reuse to accelerate VLA inference while preserving action accuracy.
+Figure 2: VLA-Cache accelerates the VLA’s language decoding process across timesteps via the following two steps: (a) Dynamic Token Selection reuses static tokens across frames while preserving task-relevant ones; (b) Adaptive Token Caching dynamically adjusts reuse ratios per decoder layer based on attention patterns.
+3.2 Temporal Redundancy in Robotic Perception
+In closed-loop robotic manipulation, consecutive visual frames often share large portions of static content.
+As illustrated in Figure 2, background regions and stationary objects typically exhibit negligible changes between adjacent frames.
+However, most existing Vision-Language-Action (VLA) models discard visual representations after each timestep and recompute all visual tokens from scratch, resulting in substantial redundancy and increased inference latency.
+To address this inefficiency, we propose selectively reusing visual tokens that remain static across timesteps.
+Specifically, we identify and cache the representations of image regions with minimal visual change, allowing their Key-Value (KV) representations to be reused in the next frame.
+This strategy significantly reduces redundant computation in the VLA visual stream while preserving model performance.
 Static Token Selection.
-To enhance inference efficiency, we identify visually static tokens across consecutive frames and reuse their associated Key-Value (KV) representations.
-This process involves three steps: partitioning the input image into patches, computing patch-wise similarity, and selecting the most static patches for reuse.
-Given an image $I$ of size $H\times W$, we divide it into $N\times N$ non-overlapping patches of size $p\times p$, yielding $N^{2}$ patches $\mathcal{P}=\{\mathbf{P}_{i,j}\}$.
-Each patch $\mathbf{P}_{i,j}$ is directly represented by its raw pixel values, without additional feature encoding.
-For two consecutive frames $I_{t}$ and $I_{t-1}$, we calculate the cosine similarity between their corresponding patches:
-$\text{Sim}\big{(}\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j}\big{)}=\frac{\mathbf{P}_{t}^{i,j}\cdot\mathbf{P}_{t-1}^{i,j}}{\|\mathbf{P}_{t}^{i,j}\|_{2}\,\|\mathbf{P}_{t-1}^{i,j}\|_{2}}.$ (6)
-A patch is considered static if its similarity score exceeds a predefined threshold $\tau$. To further refine the selection, we apply a Top-$k$ filter:
-$\mathcal{P}_{\mathrm{top\text{-}k}}=\mathrm{Top}\text{-}k\Big{(}\{\mathbf{P}_{t}^{i,j}\mid\mathrm{Sim}(\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j})\geq\tau\}\Big{)}.$ (7)
-Thus, patches with minimal visual variation are identified for reuse, while those displaying notable differences are recomputed.
-This approach accurately selects truly static tokens across consecutive frames, significantly reducing redundant computations and accelerating inference without compromising overall performance.
+Given an image $I\in\mathbb{R}^{H\times W\times 3}$, we divide it into $N\times N$ non-overlapping patches of size $p\times p$, yielding a set of raw pixel patches $\mathcal{P}_{t}=\{\mathbf{P}_{t}^{i,j}\}$.
+For each patch $\mathbf{P}_{t}^{i,j}$ in the current frame and its corresponding patch $\mathbf{P}_{t-1}^{i,j}$ in the previous frame, we compute cosine similarity:
+$\text{Sim}\big(\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j}\big)=\frac{\mathbf{P}_{t}^{i,j}\cdot\mathbf{P}_{t-1}^{i,j}}{\|\mathbf{P}_{t}^{i,j}\|_{2}\cdot\|\mathbf{P}_{t-1}^{i,j}\|_{2}}.$ (4)
+A patch is considered visually static if its similarity exceeds a threshold $\tau$. We further apply a Top-$k$ filter to retain the most stable tokens:
+$\mathcal{P}_{\mathrm{static}}=\mathrm{Top}\text{-}k\Big(\{\mathbf{P}_{t}^{i,j}\mid\mathrm{Sim}(\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j})\geq\tau\}\Big).$ (5)
+This simple yet effective approach accurately selects truly static tokens across consecutive frames, significantly reducing redundant computations and accelerating inference without compromising overall performance.
+Table 1: Comparison of VLA-Cache’s core token selection strategies on OpenVLA using the LIBERO Spatial benchmark.
+Method SR (%) $\uparrow$ Latency (ms) $\downarrow$
+OpenVLA 84.4 51.56
++ Static Token 74.2 31.03
++ Evict Task-Relevant 82.6 31.03
++ Layer Adaptive 83.8 32.22
+3.3 Retaining Task-Relevant Information
+While visual similarity offers a practical signal for reusing static regions, not all visually static tokens are safe to reuse.
+In robotic control tasks, certain regions, such as the gripper or target object, though visually unchanged, are semantically dynamic and critical for precise action generation.
+Directly reusing all visually static tokens, without considering their semantic role, can lead to serious performance degradation.
+As shown in Table 1, naive static token reuse lowers the success rate from the OpenVLA baseline of 84.4% to just 74.2%.
+This degradation occurs because task-relevant tokens, though minimally changed at the pixel level, exhibit greater feature sensitivity to subtle environmental changes.
+Unlike vision-language models (VLMs), VLA models must track object states and interactions over time, making them more dependent on accurate visual encoding.
+Therefore, to ensure alignment with the latest environment state, task-relevant regions must be recomputed each step.
 Evicting Task-Relevant Tokens.
-While many tokens remain visually static across consecutive frames, certain tokens that are highly relevant to the task (e.g., the gripper or the target object) cannot be directly reused.
-In our experiments, we observed that as the robotic arm approaches the target object, small but critical changes in the target region often occur.
-These changes are primarily caused by factors such as lighting variations or camera noise.
-Recomputing tokens around the target region before the gripper reaches it significantly improves task accuracy.
-To address this issue, we leverage the attention mechanism of the VLA model to identify task-relevant tokens in the next frame that require recalculation.
-Specifically, we calculate the text-to-vision attention scores from the cross-attention module of the VLA model to determine the task relevance of each vision token.
-For each layer $l$, the attention weights are represented as $\mathbf{A}^{l}\in\mathbb{R}^{N_{\text{heads}}\times N_{\text{tokens}}\times N_{\text{tokens}}}$.
-From these weights, we extract the attention mapping from text tokens to vision tokens:
+To avoid reusing semantically critical but visually static tokens, we propose a lightweight filtering mechanism using cross-attention scores from the language decoder.
+For each decoder layer $l$, we extract the text-to-vision attention matrix $\mathbf{A}^{l}_{\text{vis-text}}$ from the full attention tensor $\mathbf{A}^{l}\in\mathbb{R}^{N_{\text{heads}}\times N_{\text{tokens}}\times N_{\text{tokens}}}$ as:
 $\mathbf{A}^{l}_{\text{vis-text}}=\mathbf{A}^{l}[:,\,v_{\text{start}}:v_{\text{end}},\,t_{\text{start}}:t_{\text{end}}],$ (6)
@@
 Tokens that are both static and highly task-relevant are removed from the reusable token set to ensure they are recomputed in the current step:
-$\mathcal{P}_{\mathrm{final}}=\mathcal{P}_{\mathrm{static}}\;\setminus\;\mathcal{P}_{\text{task-relevant}}.$ (10)
-By filtering out high-attention tokens, this approach ensures that the VLA model accurately updates the critical objects in the scene, thereby preventing failures caused by reusing outdated static tokens.
-This method effectively balances computational efficiency with task precision, enabling robust performance in robotic control tasks.
+$\mathcal{P}_{\mathrm{reuse}}=\mathcal{P}_{\mathrm{static}}\;\setminus\;\mathcal{P}_{\text{task-relevant}}.$ (8)
+By filtering out semantically significant tokens from the static reuse set, our method restores the degraded success rate from 74.2% to 82.6%, while maintaining the computational gains in FLOPs and latency.
+This balance between task fidelity and efficiency illustrates the value of cross-modal attention as a lightweight signal for safe token reuse in VLA models.
+Algorithm 1 Dynamic Token Selection
+1:Input: Frames $\{I_{t-1},I_{t}\}$, thresholds $\tau$, $\tau_{\text{task}}$, hyperparameter $k$
+2:Output: Reusable token indices $\mathcal{P}_{\mathrm{final}}$
+[…]
+4: Patchify both frames $I_{t-1},I_{t}$ and compute $\text{Sim}(\mathbf{P}_{t},\mathbf{P}_{t-1})$ among related patches
+5: Select $\mathcal{P}_{\text{static}}$ where similarity $\geq\tau$
+6: Apply top-$k$ filtering to refine static token selection
+[…]
+8: Compute text-to-vision attention scores $\mathbf{S}_{\text{task-relevance}}$
+9: Select $\mathcal{P}_{\text{task-relevant}}$ where attention $\geq\tau_{\text{task}}$
+10: Compute reusable tokens: $\mathcal{P}_{\mathrm{final}}=\mathcal{P}_{\text{static}}\setminus\mathcal{P}_{\text{task-relevant}}$
+11:return $\mathcal{P}_{\mathrm{final}}$
+Algorithm 2 Adaptive Token Caching
+1:Input: Token indices $\mathcal{P}_{\mathrm{final}}$, previous KV cache $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$, representations $\mathbf{H}_{t}^{l}$
+2:Output: Updated KV cache $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
+3: Compute entropy $R_{\text{cum}}^{l}$, reuse ratio $\alpha^{l}$, subset $\mathcal{P}_{\mathrm{reuse}}\subseteq\mathcal{P}_{\mathrm{final}}$
+4:for each layer $l$ and token $i$ do
+5:  if $i\in\mathcal{P}_{\mathrm{reuse}}$ then
+6:   Reuse cached values: $\mathbf{K}_{t}^{l}(i)=\mathbf{K}_{t-1}^{l}(i),\mathbf{V}_{t}^{l}(i)=\mathbf{V}_{t-1}^{l}(i)$
+7:  else
+[…]
+9:  end if
+10:end for
+11:return $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
 3.4 Layer Adaptive Token Reusing
 While static token selection and task-relevance filtering eliminate a large portion of redundant computation, we observe that attention distributions within the VLA decoder vary significantly across different layers.
-This finding is consistent with observations reported by prior work FastV (Chen et al., 2025), indicating that both VLA and VLM decoders exhibit similar patterns of attention flow: early layers display dispersed attention, followed by fluctuations in intermediate layers, and eventually a partial rebound near the final layers.
+This finding is consistent with observations reported by prior work FastV [29], indicating that both VLA and VLM decoders exhibit similar patterns of attention flow: early layers display dispersed attention, followed by fluctuations in intermediate layers, and eventually a partial rebound near the final layers.
 To account for these differences, we propose a layer-adaptive strategy that adjusts the fraction of reused tokens based on each layer’s attention concentration.
-Specifically, we quantify the attention distribution at layer $l$ via an entropy measure, following the same mean-attention computation described in equation  8.
-Let $H^{l}$ denote the resulting entropy.
-We then define an entropy ratio $R^{l}$, which captures how much more concentrated the attention is in layer $l$ compared to layer $l-1$:
-$R^{l}\;=\;\frac{H^{l-1}-H^{l}}{H^{l-1}}.$ (11)
-A positive $R^{l}$ indicates that the attention distribution at layer $l$ is more focused than that of layer $l-1$.
-We accumulate these ratios across layers to obtain a cumulative score $R_{\text{cum}}^{l}=\sum_{j=1}^{l}R^{j}$, which in turn determines the proportion $\alpha^{l}$ of static tokens (from $\mathcal{P}_{\mathrm{final}}$) that are reused at layer $l$.
+Specifically, we quantify the attention distribution at layer $l$ via an entropy measure, following the same mean-attention computation described in Eq.
+6.
+Let $\mathcal{E}^{l}$ denote the resulting entropy.
+We then define an entropy ratio $R^{l}=(\mathcal{E}^{l-1}-\mathcal{E}^{l})/\mathcal{E}^{l-1}$, which captures how much more concentrated the attention is in layer $l$ compared to layer l1l-1.
+A positive $R^{l}$ indicates that the attention distribution at layer $l$ is more focused than that of layer l1l-1.
+We accumulate these ratios across layers to obtain a cumulative score, which in turn determines the proportion $\alpha^{l}$ of static tokens (from $\mathcal{P}_{\mathrm{final}}$) that are reused at layer $l$.
 Formally,
-[…]
+$\alpha^{l}\;=\;\min\Bigl(k\sum_{j=1}^{l}R^{j},\;1\Bigr),$ (9)
 where $k$ is a hyperparameter that governs the impact of attention concentration.
@@
 By selectively retaining only the most relevant tokens at each layer, our method significantly reduces redundant computations while maintaining reliable action prediction.
-Algorithm 1 Adaptive Token Caching
-1:  Input: Frames $\{I_{t-1},I_{t}\}$, previous KV cache $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$, thresholds $\tau,\tau_{\text{task}}$, hyperparameter $k$
-2:  Output: Updated KV cache $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
-3:  Static Token Selection:
-4:  Patchify $I_{t-1},I_{t}$ and compute $\text{Sim}(\mathbf{P}_{t},\mathbf{P}_{t-1})$
-5:  Select $\mathcal{P}_{\text{static}}$ where similarity $\geq\tau$
-6:  Apply top-$k$ filtering to refine static token selection
-7:  Evict Task-Relevant Tokens:
-8:  Compute text-to-vision attention scores $\mathbf{S}_{\text{task-relevance}}$
-9:  Select $\mathcal{P}_{\text{task-relevant}}$ where attention $\geq\tau_{\text{task}}$
-10:  Compute reusable tokens: $\mathcal{P}_{\mathrm{final}}=\mathcal{P}_{\text{static}}\setminus\mathcal{P}_{\text{task-relevant}}$
-11:  Layer-Adaptive Token Reuse:
-12:  Compute entropy reduction $R^{l}$, $R_{\text{cum}}^{l}$; Determine reuse ratio $\alpha^{l}$; Compute reusable subset $\mathcal{P}_{\mathrm{reuse}}\subseteq\mathcal{P}_{\mathrm{final}}$
-13:  for each layer $l$ and token $i$ do
-14:     if $i\in\mathcal{P}_{\mathrm{reuse}}$ then
-15:        Reuse cached values:    $\mathbf{K}_{t}^{l}(i)=\mathbf{K}_{t-1}^{l}(I),\mathbf{V}_{t}^{l}(i)=\mathbf{V}_{t-1}^{l}(i)$
-16:     else
-17:        Recompute:    $\mathbf{K}_{t}^{l}(i)=W_{K}^{l}\mathbf{H}_{t}^{l}(i),\mathbf{V}_{t}^{l}(i)=W_{V}^{l}\mathbf{H}_{t}^{l}(i)$
-18:     end if
-19:     Compute self-attention:
-20:     $\mathrm{Attn}(\mathbf{Q}_{t}^{l},\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l})=\mathrm{Softmax}\!\Bigl{(}\frac{\mathbf{Q}_{t}^{l}(\mathbf{K}_{t}^{l})^{\top}}{\sqrt{D}}\Bigr{)}\mathbf{V}_{t}^{l}$
-21:  end for
-22:  return Updated KV cache $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$
-Inference with Cached Representation.
-During inference, VLA-Cache selectively caches and reuses previously computed representations to reduce redundant computations while maintaining accurate robotic action prediction.
-Instead of recomputing all tokens at each timestep, the model distinguishes between static and dynamic tokens.
-Static tokens, which exhibit minimal visual change across consecutive frames, retain their cached key-value (KV) representations from the previous timestep, bypassing unnecessary computation.
-Dynamic tokens, which undergo significant changes or are task-relevant, are recomputed to ensure precise motion execution.
-This selective token reuse is performed at each layer, guided by entropy-based attention analysis to dynamically adjust reuse ratios.
-The full inference procedure, including static token selection, task-relevance filtering, and layer-adaptive token reuse, is outlined in Algorithm 1, with further details provided in Appendix A.2.
+4 Implementations
+4.1 Cross-Frame Visual Token Caching
+During inference, VLA-Cache accelerates robotic action prediction by reusing previously computed key-value (KV) representations of visual tokens across time.
+Instead of recomputing all visual tokens in each time step, the model identifies a subset of static tokens, those that exhibit minimal change across frames, and reuses their cached representations from the previous time step.
+In contrast, dynamic tokens, which undergo significant visual change or are task-relevant, are freshly computed to maintain accurate action generation.
+At each timestep $t$, given the visual token sequence $\mathbf{H}_{t}$, the model identifies a subset $\mathcal{P}_{\text{reuse}}$ of tokens that remain unchanged from the previous frame and reuses their cached representations:
+$\mathbf{K}_{t}(i)=\begin{cases}\mathbf{K}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}}\\ W_{K}\mathbf{H}_{t}(i),&\text{otherwise}\end{cases},\quad\mathbf{V}_{t}(i)=\begin{cases}\mathbf{V}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}}\\ W_{V}\mathbf{H}_{t}(i),&\text{otherwise}\end{cases}.$ (10)
+This design avoids redundant computation for static tokens while ensuring dynamic or task-relevant inputs are freshly computed.
+Most existing approaches reduce computation by pruning or merging tokens within a single frame [29, 30, 31, 32, 33].
+In contrast, VLA-Cache exploits temporal redundancy by caching and reusing visual tokens across frames, making it better aligned with the closed-loop nature of robotic control.
+Notably, it is compatible with high-frequency architectures and directly alleviates the decoding bottleneck.
+Since it requires no model modification or retraining, VLA-Cache serves as a plug-and-play optimization for efficient robotic inference.
+An overview of the inference procedure is shown in Algorithm 1 and Algorithm 2, detailing token selection and adaptive caching.
+Additional implementation details are provided in Appendix D.
 4.2 Theoretical Analysis of Computational Complexity
-Computational Cost Reduction. In standard VLA inference, each Transformer layer processes $L$ tokens, with a total FLOP cost per layer:
-$\text{FLOPs}\approx 4LD^{2}+2L^{2}D+2LDM.$ (13)
-Our method reduces the effective token count per layer to $L_{r}=\alpha\times\mathcal{P}_{\mathrm{final}}$, leading to theoretical savings:
-$\Delta\text{FLOPs}_{\mathrm{layer}}\approx 4L_{r}D^{2}+2L_{r}^{2}D+2L_{r}DM.$ (14)
-These savings scale across multiple layers, significantly lowering computation while maintaining accuracy.
 Overhead of Token Selection.
@@
 The entropy-based layer-adaptive strategy incurs an additional $\mathcal{O}(L^{2}D)$ complexity, which remains significantly lower than the baseline per-layer cost.
+Computational Cost Reduction. In standard VLA inference, each Transformer layer processes $L$ tokens, with a total FLOP cost per layer:
+$\text{FLOPs}\approx 4LD^{2}+2L^{2}D+2LDM.$ (11)
+VLA-Cache reduces effective token count per layer to $L_{r}=\alpha\times\mathcal{P}_{\mathrm{final}}$, leading to theoretical savings:
+$\Delta\text{FLOPs}_{\mathrm{layer}}\approx 4L_{r}D^{2}+2L_{r}^{2}D+2L_{r}DM.$ (12)
 Total Complexity Reduction. Bringing all components together, the theoretical overall FLOP reduction per layer is:
-$\Delta\text{FLOPs}_{\text{total}}\approx$ $\Bigl{(}4L_{r}D^{2}+2L_{r}^{2}D+2L_{r}DM\Bigr{)}$ (15)
-$-\Bigl{(}H^{2}+L_{t}L_{v}D+L^{2}D\Bigr{)}.$
-Further Derivations and Complexity Details. See Appendix A.1 for detailed derivations, including static token selection costs, attention filtering complexity, and layer-adaptive entropy calculations.
+$\Delta\text{FLOPs}_{\text{total}}\approx\Bigl(4L_{r}D^{2}+2L_{r}^{2}D+2L_{r}DM\Bigr)-\Bigl(H^{2}+L_{t}L_{v}D+L^{2}D\Bigr).$ (13)
+Please refer to Appendix C for detailed derivations, including static token selection costs, attention filtering complexity, and layer-adaptive entropy calculations.
 5 Experiment
-5.1 Experimental Settings
-To validate the generalization of VLA-Cache, we implement our method in both simulation and real-world settings.
-In simulation, we evaluate VLA-Cache on two open-source VLA models: OpenVLA (Kim et al., 2024) and CogAct (Li et al., 2024a), using the LIBERO benchmark (Liu et al., 2024a) and SIMPLER environment (Li et al., 2024b), respectively.
-For fair comparison, we adopt two token-pruning acceleration methods, SparseVLM (Zhang et al., 2024) and FastV (Chen et al., 2025), as baselines in the LIBERO environment.
-OpenVLA processes 256 visual tokens per image, and all comparative experiments use an identical token pruning and reuse count, setting Top-K to 100 tokens.
-For VLA-Cache, the default parameters are set as follows: similarity score threshold $\tau=0.996$, top-$k=100$, and task relevance threshold $\tau_{task}=0.5$.
-Additionally, we use OpenVLA’s default caching mechanism and other default parameters for inference acceleration.
-In the SIMPLER environment, we apply VLA-Cache with the same parameter settings as OpenVLA while using CogAct’s default hyperparameters.
-For real-world evaluation, we fine-tune OpenVLA for deployment on a Kinova Jaco2 6-DoF manipulator.
-The model is trained for 50,000 steps with default parameters.
-During testing, we modify only the similarity score threshold to $\tau=0.85$, keeping all other parameters unchanged.
+To validate the effectiveness of VLA-Cache, we evaluate our method in both simulation and real-world settings.
+In simulation, we evaluate VLA-Cache on three open-source VLA models: OpenVLA [11], OpenVLA-OFT [20] and CogAct [19], using the LIBERO benchmark [17] and SIMPLER environment [18], respectively.
 All experiments are conducted on an NVIDIA RTX 4090 GPU.
-5.1.1 LIBERO Benchmark
-The LIBERO Benchmark (Liu et al., 2024a) is designed to evaluate lifelong robotic manipulation across four diverse task suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long.
-Each suite introduces distinct variations in spatial layouts, object selection, task objectives, or long-horizon planning, challenging models to generalize across different manipulation scenarios.
-Following the experimental setup of OpenVLA, we conduct evaluations on all four LIBERO suites, each containing ten subtasks.
-To ensure consistency, we use the same machine and official fine-tuned weights.
-For each subtask, multiple evaluation episodes are run, measuring key performance metrics such as success rate and inference latency.
-We evaluate VLA-Cache’s acceleration capabilities within the OpenVLA framework and conduct an ablation study on the top-$k$ parameter in the LIBERO environment.
-Figure 3: Simulation Tasks on LIBERO Benchmark and the SIMPLER Environment.
-Figure 4: Robot Manipulation Tasks on Jaco2 Real-world Platform.
-Figure 5: Heatmap Visualization of Robot Actions in a Simulated Environment with Prompt Descriptions and Outputs
-[…]
-LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
-Baseline (OpenVLA) 84.4% 86.6% 75.6% 53.2% 75.0% 1.864 51.91
-SparseVLM 79.8% 67.0% 72.6% 39.4% 64.7% 1.407 83.39
-FastV 83.4% 84.0% 74.2% 51.6% 73.3% 1.864 53.28
-VLA-Cache 83.8% 85.8% 76.4% 52.8% 74.7% 1.355 31.83
+5.1 Experiment Setup
+Compared Methods.
+We leverage the architectural similarity between VLA and VLM models, which allows direct application of existing VLM acceleration methods to VLA inference.
+Specifically, we adopt two state-of-the-art token-level acceleration techniques SparseVLM [30] and FastV [29] on OpenVLA as compared methods in the LIBERO benchmark.
+Evaluation Metrics.
+We evaluate VLA-Cache using four metrics: success rate, control frequency, FLOPs, and CUDA latency.
+Success rate and control frequency respectively assess task performance and the responsiveness of action prediction in closed-loop control.
+FLOPs measure theoretical computation, while CUDA latency captures actual GPU runtime.
+These two efficiency metrics are widely adopted in VLM/VLA acceleration methods.
+5.2 Evaluation Benchmark
+LIBERO.
+The LIBERO Benchmark [17] covers four task suites: Spatial, Object, Goal, and Long, each testing a different aspect of manipulation generalization.
+We follow the standard setup from OpenVLA [11] and OpenVLA-OFT [20], using official weights and machines for consistency.
+Each suite includes ten subtasks evaluated over multiple episodes.
+Figure 3: Tasks on LIBERO Benchmark, the SIMPLER Environment and Real World.
+SIMPLER.
+The SIMPLER simulator [18] offers two settings, Visual Matching and Variant Aggregation, designed to bridge simulation-to-reality gaps.
+Following CogAct’s setup [19], we evaluate both settings on a Google robot arm across four manipulation tasks.
+CogAct, which integrates a Diffusion Policy for continuous control, serves as our baseline.
+These evaluations demonstrate the generality of VLA-Cache under different action heads and simulation variations.
+Real Robot Evaluation.
+We deploy VLA-Cache on a Kinova Jaco2 manipulator equipped with a front-facing camera.
+The robot is evaluated on four tasks: PickPot, PlaceCube, PutSausage, and WipeTable, with the last task including diverse distractor objects to test robustness.
+Demonstrations are collected via teleoperation at 10 Hz using an Xbox controller, resulting in 150-200 trajectories per task.
+We fine-tune OpenVLA with LoRA [38] and evaluate on the same tasks, using the LIBERO tuning setup for consistency.
+More details about real-world experiments are available in the AppendixE.4.
+5.3 Results on Simulation Environment
 Table 2: Comparison of different VLA acceleration methods on the LIBERO benchmark.
 […]
+Spatial Object Goal Long Average
+OpenVLA 84.4% 86.6% 75.6% 53.2% 75.0% 1.864 51.91 4.23
++ SparseVLM 79.8% 67.0% 72.6% 39.4% 64.7% 1.407 83.39 3.72
++ FastV 83.4% 84.0% 74.2% 51.6% 73.3% 1.864 53.28 4.19
++ VLA-Cache 83.8% 85.8% 76.4% 52.8% 74.7% 1.355 31.83 4.59
+OpenVLA-OFT 97.8% 97.6% 97.6% 94.2% 96.8% 4.013 79.05 65.10
++ VLA-Cache 98.3% 97.5% 98.3% 95.4% 97.4% 3.097 62.59 78.98
+Table 3: Comparison of VLA-Cache within the CogACT model in the SIMPLER environment.
+[…]
 PickCan MoveNear Drawer DrawerApple Average
-Visual Matching Baseline (CogACT) 91.3% 85.0% 71.8% 50.9% 74.8% 1.847 54.29
-VLA-Cache 92.0% 83.3% 70.5% 51.6% 74.4% 1.496 39.63
-Variant Aggregation Baseline (CogACT) 89.6% 80.8% 28.3% 46.6% 61.3% 1.807 53.54
-VLA-Cache 91.7% 79.3% 32.5% 45.8% 62.3% 1.493 39.11
-Table 2: Comparison of VLA Cache in CogACT model within the SIMPLER environment.
-5.1.2 SIMPLER Environments
-The SIMPLER simulation environment (Li et al., 2024b) is designed to closely mirror real-world robotic setups, facilitating more realistic evaluations.
-SIMPLER supports two evaluation settings: Visual Matching (aiming to minimize visual discrepancies between simulation and reality) and Variant Aggregation (building on Visual Matching by introducing variations in lighting, background, table textures, etc.).
-We adopt the same SIMPLER configuration used by CogAct and evaluate two settings, Visual Matching and Variant Aggregation, on a Google robot arm.
-Each setting contains four tasks: 1) Pick coke can(PickCan), 2) Move near(MoveNear), 3) Open/close drawer(Drawer), 4) Open top drawer and place apple(DrawerApple).
-To validate the general applicability of our method, we use CogAct as the baseline model within the SIMPLER environment.
-CogAct extends OpenVLA by incorporating a Diffusion Policy head to predict actions in continuous space, a common strategy among recent VLA models.
-Similar to OpenVLA, CogAct provides publicly available weights and code for reproducible simulation-based evaluations.
-Through these experiments, we try to verify its potential as a general VLA acceleration technique.
-5.1.3 Real-world Robot Setting
-We further evaluate our method on a Kinova Jaco2 6-DoF manipulator, equipped with a Sony AX53 camera positioned in front of the arm to capture the manipulation scene.
-Tasks.
-We design four real-world robotic manipulation tasks:1) Pick up the orange pot (PickPot),2) Place the blue cube in the box (PlaceCube), 3) Put the sausage in the blue pan (PutSausage),4) Wipe the table (WipeTable).
-For the wiping task, we place objects of various shapes and colors on the table to challenge the model’s robustness to unseen objects.
-Data Collection.
-Using an Xbox Gamepad, we teleoperate the Jaco2 arm and collect demonstrations at a frequency of 10 Hz, recording RGB images and robot states throughout each trajectory.
-We gather 150–200 demonstrations per task to form the training dataset.
-Baseline.
-We fine-tune OpenVLA via LoRA (Low-Rank Adaptation (Hu et al., 2021)) on each task and then deploy it on the Jaco2 to assess real-world performance.
-The fine-tuning process follows the same hyperparameter setup as in LIBERO.
-5.2 Results on Simulation Environment
+Matching CogACT 91.3% 85.0% 71.8% 50.9% 74.8% 1.847 54.29 12.42
+     + VLA-Cache 92.0% 83.3% 70.5% 51.6% 74.4% 1.496 39.63 14.66
+Aggregation CogACT 89.6% 80.8% 28.3% 46.6% 61.3% 1.807 53.54 12.36
+     + VLA-Cache 91.7% 79.3% 32.5% 45.8% 62.3% 1.493 39.11 14.48
+Main Results on LIBERO.
 […]
-PickPot PlaceCube PutSausage WipeTable Average
-Baseline (OpenVLA) 95.0% 83.3% 80.0% 70.0% 82.1% 1.814 64.16
-VLA-Cache 90.0% 90.0% 85.0% 73.0% 84.5% 1.303 51.85
-Table 3: Comparison of success rate on real robot tasks.
-[…]
-Static Token 74.2% 1.304 31.03
-Task Relevant 82.6% 1.304 31.03
-Layer Adaptive (Ours) 83.8% 1.382 32.22
-Table 4: Comparison of different token selection methods.
-  #Tokens Methods SR % $\uparrow$ FLOPs $\downarrow$ Time(ms) $\downarrow$
+Table 4: Ablation on token pruning/reuse in LIBERO-Spatial using OpenVLA (256 vision tokens). Best values are in bold.
+#Tokens Methods SR % $\uparrow$ FLOPs $\downarrow$ Latency (ms) $\downarrow$
 0 Baseline 84.4 1.888 52.37
@@
 Ours 68.3 0.823 30.29
-Table 5: LIBERO-Spatial benchmark comparing different methods based on pruning or reuse tokens. Best values in each column are highlighted in bold.
-Main Results on LIBERO.
-Table 1 summarizes the success rate, FLOPs, and CUDA inference time across the four LIBERO task suites.
-Compared to the baseline, VLA-Cache achieves a 27.31% FLOPs reduction and a 1.63$\times$ speedup, with only a minimal 0.3% drop in the overall success rate, confirming its efficiency without compromising accuracy.
-While VLA-Cache slightly underperforms on spatial, object, and long-horizon tasks, it notably surpasses the baseline on the goal suite, highlighting its robust adaptability.
-Notably, both FastV and SparseVLM fail to improve inference time, often exceeding the baseline.
-FastV merely masks tokens in attention computation without reducing GPU workload, introducing additional masking overhead, while SparseVLM incurs extra cost from token pruning, merging, and recycling.
-These methods were originally designed for long-sequence VLM tasks, where efficiency gains are more pronounced, but struggle in VLA settings, where action sequences are typically short.
-In contrast, VLA-Cache leverages inter-frame token reuse with negligible overhead, making it effective for robotic action generation.
-Ablation on Reusing/Pruning Rate.
-Table 5 presents a comparative analysis of different methods through an ablation study on the top-$k$ parameter, while also varying the number of pruned or reused tokens.
-For all methods, increasing the number of reused/pruned tokens leads to a decline in success rates, highlighting the importance of retaining sufficient information for accurate action predictions.
-Our approach maintains a relatively stable success rate at a reuse rate of 40%, but a more noticeable drop occurs when reusing 200 tokens.
-In contrast, FastV removes a significant portion of tokens entirely, and its performance drop primarily results from the loss of critical visual details rather than potential errors in token reuse.
-Additionally, SparseVLM and FastV incur longer inference times due to extra attention operations during GPU execution, whereas VLA-Cache directly updates KV cache entries, making inference more efficient.
-Token Selection Strategies As presented in Table 4, we evaluate different token selection strategies for VLA-Cache.
-The direct reuse of static tokens results in a significant decline in success rate.
-In contrast, filtering task-relevant tokens maintains a success rate of 82.6%.
-Notably, our layer-adaptive approach further enhances the success rate to 83.8%, albeit with a marginal increase in computational overhead from 1.304 FLOPs to 1.382 FLOPs.
-This overhead arises because fewer tokens are reused in the early layers, which ensures the retention of critical updates in deeper layers, thereby improving overall performance.
+As illustrated in Figure 4, VLA-Cache effectively reduces visual computation redundancy during robotic manipulation by precisely identifying static tokens and filtering out task-relevant regions in real time.
+Notably, OpenVLA-OFT takes both fixed third-person and dynamic wrist-camera views as inputs.
+Its strong performance demonstrates that VLA-Cache not only enhances control frequency but also maintains robustness under dynamic viewpoints and camera shifts.
+Ablation on Token Reusing/Pruning Rate.
+Table 4 presents an ablation study varying the number of reused/pruned tokens.
+For all methods, aggressive token reduction harms success rate, underscoring the need to preserve informative content.
+VLA-Cache maintains stable performance at moderate reuse rates (i.e., 100 tokens), while FastV and SparseVLM suffer larger drops due to loss of critical visual details.
+By directly updating KV entries, VLA-Cache remains more efficient and robust under different token reuse configurations.
+Token Selection Strategies.
+As shown in Table 1, directly reusing all static tokens leads to a notable drop in success rate 74.2%, indicating that visual similarity alone is insufficient for reliable reuse in robotic control.
+By filtering out task-relevant tokens based on decoder attention, our method recovers performance to 82.6%.
+Introducing a layer-adaptive strategy further improves accuracy to 83.8%, with minimal increase in CUDA latency.
 Main Results on SIMPLER
-Table 2 indicates that the VLA-Cache exhibits comparable success rates than the baseline CogACT in the SIMPLER environment while substantially reducing computational overhead.
-Specifically, VLA-Cache achieves an average success rate of 74.4% in the standard Visual Matching setting compared to 74.8% for CogACT and 62.3% in the Variant Aggregation setting versus 61.3% for CogACT.
-The efficiency gains are evident in the FLOPs and inference time measurements.
-VLA-Cache achieves roughly 20% fewer FLOPs than the baseline, coupled with a 1.37× reduction in inference latency.
-Notably, these results highlight the portability of VLA-Cache across different action heads, establishing it as a general acceleration strategy for VLA.
+Table 3 indicates that VLA-Cache exhibits success rates comparable to the CogACT baseline in the SIMPLER environment while substantially reducing computational overhead.
+[…]
+Figure 4: Visualization of VLA-Cache token reuse across settings.
+(a) LIBERO simulation with OpenVLA.
+(b) Real-world task under dynamic background.
+(c) and (d) Main and wrist camera views from OpenVLA-OFT.
+Blue: static tokens, Yellow: task-relevant, Red: overlapping.
+VLA-Cache reduces redundant computation and preserves accuracy under varying conditions.
+Table 5: Comparison of success rate on real robot tasks.
+[…]
+PickPot PlaceCube PutSausage WipeTable Average
+OpenVLA 95.0% 83.3% 80.0% 70.0% 82.1% 1.814 64.16 4.02
+   + VLA-Cache 90.0% 90.0% 85.0% 73.3% 84.6% 1.303 51.85 4.21
 5.4 Results on Real Robot
 Table 5 illustrates the performance of VLA-Cache in real-world robotic tasks.
-Among the four tasks, PickPot shows a slightly lower success rate than the baseline, whereas VLA-Cache exceeds the baseline on the other three tasks.
+Among the four tasks, PickPot shows a slightly lower success rate than the baseline, whereas VLA-Cache exceeds the baseline on other three tasks.
 The method also achieves considerable reductions in FLOPs and inference time.
-Overall, VLA-Cache improves the average success rate by 2.4%, an outcome that, although unexpected, is potentially explained by a more robust baseline model, which we trained extensively using additional data and more training steps than the LIBERO fine-tuning.
+Overall, VLA-Cache improves the average success rate by 2.4%, likely due to reduced interference from redundant visual tokens and enhanced decision robustness.
 With improved robustness as a foundation, VLA-Cache’s ability to prune or reuse redundant tokens may further enhance the model’s resilience, thus yielding higher success rates.
+Performance under Dynamic Background.
+To assess robustness, we introduced background motion (e.g., human hands and moving objects) in the PickPot task.
+As shown in Table 7, success rate of baseline dropped from 95% to 80% under noise.
+With VLA-Cache, the same success rate was maintained while reducing FLOPs by 42% and latency by 35%.
+These results highlight VLA-Cache’s ability to filter out transient or irrelevant tokens, preserving both efficiency and stability in dynamic real-world settings.
+Figure 4 visualizes this effect, showing robust token selection despite background disturbances.
 6 Conclusion
-In this paper, we introduce VLA-Cache, a training-free method for VLA that selectively reuses static tokens while filtering out task-relevant ones, reducing redundant computation without sacrificing accuracy.
-Additionally, our layer-adaptive token reuse strategy improves model success rates by adjusting token reuse based on attention concentration.
-Extensive experiments on two VLA models, OpenVLA and CogAct, across two simulation environments, LIBERO and SIMPLER, demonstrate that VLA-Cache achieves a 1.7$\times$ speedup while maintaining performance.
-Appendix A Appendix
-A.1 Complexity Analysis Details
+[…]
+Acknowledgments
+This work was supported in part by the Australian Research Council under Projects DP240101848 and FT230100549.
+We would also like to thank the anonymous reviewers and area chair for their constructive feedback, which helped improve the clarity and quality of this paper.
+Appendix A Limitations
+In this section, we discuss the potential limitations of the proposed VLA-Cache: (i) In dynamic environments with substantial background or object motion, the number of non-reusable tokens increases, reducing acceleration gains.
+As visualized in Figure 4, dynamic regions (highlighted in yellow) require full recomputation.
+(ii) Our experiments focus on three state-of-the-art open-source VLA architectures (OpenVLA [11], CogAct [19], OpenVLA-OFT [20]) based on LLaMA2 [15] decoders.
+The applicability of VLA-Cache to emerging VLA systems with different backbones (e.g., Gemma2 [16] in $\pi_{0}$ [24]) or more complex VLA systems remains an open direction for future work.
+Appendix B Impact Statement
+While VLA-Cache demonstrates robust performance even under dynamic conditions, we emphasize the importance of careful deployment and continuous monitoring to ensure safe, interpretable, and reliable behavior in real-world robotic systems.
+Appendix C Complexity Analysis Details
 Static Token Selection. We compute patch-wise similarity for visual tokens:
@@
 This confirms that dynamic token reuse significantly reduces computation without sacrificing model performance.
-A.2 VLA Cache Inference
-Having identified both static tokens and task-relevant tokens as described in Section 4.1, we now detail how these tokens are reused in the VLA decoder via KV caching.
-Below, we first present a high-level overview of the procedure, then introduce several formulas to clarify the partial update of key-value (KV) representations under our selective token reuse strategy.
-Overall Procedure.
-At time step $t-1$, while predicting the action for frame $I_{t-1}$, the model produces the attention weights (used to evaluate task relevance) and the KV cache, $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$ for each layer $l$.
-When a new frame $I_{t}$ arrives, we compare it with $I_{t-1}$ (6) to obtain a set of static tokens sorted by similarity scores, and further filter out any high task-relevance tokens (9).
-Simultaneously, we determine a layer-specific reuse ratio $\alpha^{l}$  (12) based on the entropy of attention distributions at time $t-1$.
-This yields:
-$\mathcal{P}_{\mathrm{final}}\;\text{and}\;\{\alpha^{l}\}_{l=1}^{\Omega},$ (21)
-where $\mathcal{P}_{\mathrm{final}}$ denotes the final subset of tokens eligible for reuse, and $\Omega$ is the total number of decoder layers.
-Selective Token Reuse.
-During decoding at time $t$, each decoder layer $l$ reuses only the top-$k$ tokens (determined by $\alpha^{l}$ and similarity ranking) from $\mathcal{P}_{\mathrm{final}}$.
-Let $\mathcal{P}_{\mathrm{reuse}}\subseteq\mathcal{P}_{\mathrm{final}}$ be this chosen set of reusable tokens, with positions (IDs) $\{i\mid i\in\mathcal{P}_{\mathrm{reuse}}\}$.
-For tokens not in $\mathcal{P}_{\mathrm{reuse}}$, we recompute their key-value (KV) representations from scratch.
-Concretely, let $\mathbf{H}_{t}^{l}\in\mathbb{R}^{L\times D}$ be the hidden states at layer $l$ for the new frame $I_{t}$.
-The per-layer self-attention requires three components: queries $\mathbf{Q}_{t}^{l}$, keys $\mathbf{K}_{t}^{l}$, and values $\mathbf{V}_{t}^{l}$.
-We define:
-$\mathbf{Q}_{t}^{l}(i)\;=\;W_{Q}^{l}\,\mathbf{H}_{t}^{l}(i),$ (22)
-for all token indices $i\in\{1,\ldots,L\}$, where $W_{Q}^{l}\in\mathbb{R}^{D\times D}$ is the learnable projection of queries.
-For $\mathbf{K}_{t}^{l}$ and $\mathbf{V}_{t}^{l}$, we partially reuse cached values from time $t-1$:
-$\mathbf{K}_{t}^{l}(i)\;=\;\begin{cases}\mathbf{K}_{t-1}^{l}(i),&\text{if }i\in\mathcal{P}_{\mathrm{reuse}},\\ W_{K}^{l}\,\mathbf{H}_{t}^{l}(i),&\text{otherwise},\end{cases}$ (23)
-$\mathbf{V}_{t}^{l}(i)\;=\;\begin{cases}\mathbf{V}_{t-1}^{l}(i),&\text{if }i\in\mathcal{P}_{\mathrm{reuse}},\\ W_{V}^{l}\,\mathbf{H}_{t}^{l}(i),&\text{otherwise},\end{cases}$ (24)
-where $W_{K}^{l},W_{V}^{l}\in\mathbb{R}^{D\times D}$ are the key and value projection matrices, respectively, and $\mathbf{K}_{t-1}^{l}(i),\mathbf{V}_{t-1}^{l}(i)$ are the cached key-value vectors from the previous time step.
-This partial-update scheme allows us to skip computation for those tokens deemed both static and non-task-critical.
-Finally, the self-attention output at layer $l$ is computed as:
-$\mathrm{Attn}(\mathbf{Q}_{t}^{l},\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l})\;=\;\mathrm{Softmax}\!\Bigl{(}\tfrac{\mathbf{Q}_{t}^{l}(\mathbf{K}_{t}^{l})^{\top}}{\sqrt{D}}\Bigr{)}\;\mathbf{V}_{t}^{l}.$ (25)
-Only the newly computed tokens update $\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}$, while reused tokens retain their entries from $\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}$.
+Appendix D Inference Detail of VLA-Cache
+Inference Procedure.
+VLA-Cache accelerates robotic action prediction by reusing static visual tokens across frames during inference.
+At each timestep $t$, current visual tokens $\mathbf{H}_{t}$ are compared with those from the previous frame to identify unchanged regions.
+Tokens that are both visually static and task-irrelevant are reused via cached key-value (KV) entries, while task-relevant or dynamic tokens are recomputed.
+This selective reuse substantially reduces visual processing cost without altering model architecture or training.
+Token Reuse Mechanism.
+Our implementation modifies the VLA decoder’s forward pass as follows:
+  • Position and Attention Masking.
+We maintain a cache_position array to mark tokens requiring recomputation.
+Static tokens retain their previous position encodings, allowing attention masks to be pruned to match the reduced token set.
+  • Rotary Embedding. For recomputed tokens, rotary embeddings are applied to introduce positional information. Tokens that are skipped retain their previous encoded states.
+  • Dynamic Cache Updates.
+Newly computed tokens update their respective $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$ entries, while reused tokens inherit values from the previous frame’s cache $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$.
+Due to the permutation invariance of Transformers, this partial update yields valid attention results.
+This strategy is fully compatible with standard KV caching in autoregressive decoding.
+The largest computational gain occurs when generating the first action token at each timestep; subsequent tokens are decoded autoregressively without additional cost.
+Experimental Settings.
+All experiments are conducted using OpenVLA with 256 visual tokens.
+Unless specified otherwise, we use a static token similarity threshold $\tau=0.996$, top-$k=100$ for retained static tokens, and a task-relevance threshold $\tau_{\text{task}}=0.5$.
+These parameters are applied consistently across all simulated and real-world settings, including SIMPLER with CogAct.
+For real-world Jaco2 experiments, we slightly reduce the similarity threshold to $\tau=0.85$ to accommodate environmental noise.
+Training on the real robot used LoRA-based fine-tuning for 50,000 steps, and all evaluations were performed on an NVIDIA RTX 4090 GPU.
+Appendix E More Experiment Results
+E.1 Simulation Experiment
+LIBERO Task Definitions.
+Similarly, we also utilize all task suites provided in LIBERO for our evaluations.
+The Robosuite-based robot setup includes the following tasks: 1) “place bowl on plate with spatial variation” (e.g., drawer positions), 2) “pick object” (e.g., ketchup, bowl, apple), 3) “(open / close) target drawer; action object” (e.g., “open top drawer; place apple into drawer”), and 4) “achieve goal using shared objects” (e.g., rearranging spatial relationships or altering object states).
+SIMPLER Task Definitions.
+We utilize all task variants provided in SIMPLER for our evaluations, which include the Google robot setup with the following tasks: 1) “pick Coke can”, 2) “move obj1 near obj2”, 3) “(open / close) (top / middle / bottom) drawer”, and 4) “open top drawer; place apple into top drawer”.
+Evaluations for the Google robot setup are provided for both Visual Matching (VM) and Variant Aggregations (VA).
 Implementation Details.
-Listing A.2 outlines our partial token reuse within the VLA decoder:
-  1. 1.
-    Position IDs and Causal Masks.
-We maintain a cache_position array reflecting the positions of tokens that require computation.
-Tokens not recomputed do not receive new position IDs, enabling a pruned causal mask aligned with the updated token set.
-  2. 2.
-    Computing Rotary Embeddings.
-Once the reusable tokens are omitted from re-encoding, we apply a rotary embedding function to the reduced set of token embeddings to generate position-dependent features for the next layer.
-  3. 3.
-    KV Cache Updates.
-Only newly computed tokens update their corresponding key-value entries in $\{\mathbf{K}_{t}^{l},\mathbf{V}_{t}^{l}\}$.
-Reused tokens keep their entries from $\{\mathbf{K}_{t-1}^{l},\mathbf{V}_{t-1}^{l}\}$.
-This partial-update mechanism leverages the permutation-invariance property of Transformers, ensuring consistent attention outputs even if certain tokens are omitted from recomputation.
-Since Transformers naturally support skipping previously processed tokens in a self-attention cache, our layer-adaptive token reuse is readily integrated with standard caching strategies.
-Notably, the most significant reduction in computation occurs when generating the first action token at time $t$.
-After that, the autoregressive procedure continues as usual, incurring no additional overhead for subsequent tokens.
-Efficiency and Effect on Performance.
-By leveraging selective token reuse, VLA-Cache decreases the total number of tokens that must be re-encoded and re-attended at each time step $t$.
-This leads to a substantial reduction in decoding time, especially given the short sequences in robotic action prediction.
-Meanwhile, the complexity of predicting subsequent tokens in the same time step follows the usual KV-caching paradigm, adding minimal overhead.
-In practice, this method preserves task accuracy by forcing recomputation of highly dynamic or task-critical tokens while reusing the remainder.
-We present a detailed evaluation in Section 5, demonstrating that VLA-Cache achieves notable speedups with negligible performance degradation.
-A.3 Finetuning Data for Real Robot Experiments
+Simulated evaluations for CogACT and SIMPLER are conducted on a single NVIDIA RTX 4090 GPU in BF16 precision.
+During inference, we use DDIM sampling with 10 steps and a classifier-free guidance (CFG) coefficient of 1.5.
+Similarly, for OpenVLA and LIBERO, inference is performed on a single NVIDIA RTX 4090 GPU in BF16 precision.
+E.2 Additional Simulation Results
+Result of Subtask on LIBERO Spatial Task Suit.
+Table 6 presents detailed results on each subtask in the LIBERO-Spatial suite.
+We observe that VLA-Cache, along with methods like SparseVLM and FastV, occasionally surpasses the baseline’s success rate on individual subtasks.
+This suggests that certain redundant tokens may distract the baseline model, and pruning or reusing tokens can in fact enhance its robustness.
+Table 6: Comparison of success rates across different tasks in the LIBERO-Spatial benchmark.
+Method Success Rate % $\uparrow$ FLOPs $\downarrow$ CUDA Time (ms) $\downarrow$
+1 2 3 4 5 6 7 8 9 10 Avg
+Baseline (OpenVLA) 90 90 84 96 70 90 96 76 82 70 84.4 1.888 52.37
+SparseVLM 88 60 90 90 60 82 90 92 72 74 79.8 1.367 88.08
+FastV 92 90 90 94 58 92 90 80 78 70 83.4 1.888 54.00
+VLA-Cache 90 90 88 94 66 84 94 84 76 72 83.8 1.382 32.22
+Table 7: Real-world PickPot task under dynamic background (human/object motion).
+Method Success (%) $\uparrow$ FLOPs $\downarrow$ Latency (ms) $\downarrow$
+Baseline (OpenVLA) 95 1.800 68.03
++ Noise 80 1.807 68.22
++ Noise, VLA-Cache 80 1.275 50.59
+Visualization Results.
+Figure 5 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache in LIBERO. More visualized results are available in the supplementary materials.
+Figure 5: VLA-Cache test results and attention heat map in a simulated environment
+E.3 Additional Ablations and Comparisons
+Attention vs. object–mask proxies for task relevance.
+To validate whether attention score is a reliable proxy for task relevance, we compared our default attention–based filtering with an object-mask alternative using Efficient Track Anything[39].
+On LIBERO-Spatial with OpenVLA-OFT, attention-based proxy yields both higher success rate and lower latency than the mask variant:
+Method SR (%) $\uparrow$ FLOPs(T) $\downarrow$ Latency (ms) $\downarrow$ Freq. (Hz) $\uparrow$
+OpenVLA-OFT 97.8 3.99 78.35 65.44
+   + VLA-Cache (Attention) 98.3 3.04 61.12 81.67
+   + VLA-Cache (Object Mask) 87.4 3.15 87.49 64.78
+Table 8: Attention vs. object-mask proxies for task relevance on LIBERO-Spatial.
+While object masks provide spatial localization, they can miss fine-grained or contextual signals essential for manipulation, especially with background clutter or small, relevant parts.
+In contrast, attention scores are dynamically produced and tightly coupled with the model’s internal reasoning, providing a lightweight, task-adaptive proxy for relevance.
+Sensitivity to static-token budget $k$ and relevance threshold $\tau$.
+We further analyze the trade-off between accuracy and efficiency by varying the number of static tokens $k$ (with $\tau{=}0.5$ fixed) and the task-relevance threshold $\tau$ (with k=100k{=}100 fixed) on LIBERO-Spatial using OpenVLA-OFT.
+Results indicate robustness across a wide range; our default $(k{=}100,\ \tau{=}0.5)$ offers a strong balance.
+Method SR (%) $\uparrow$ FLOPs (T) $\downarrow$ Latency (ms) $\downarrow$ Freq. (Hz) $\uparrow$
+OpenVLA-OFT 97.8 3.995 78.35 65.44
+   + VLA-Cache (k=50k{=}50) 97.6 3.332 66.82 77.23
+   + VLA-Cache (k=80k{=}80) 97.8 3.226 66.55 78.66
+   + VLA-Cache (k=100k{=}100) 98.2 3.156 64.88 79.88
+   + VLA-Cache (k=120k{=}120) 98.0 3.109 62.99 80.56
+   + VLA-Cache (k=150k{=}150) 97.4 3.043 61.12 81.67
+   + VLA-Cache (k=180k{=}180) 96.6 2.936 60.46 82.51
+Table 9: Varying the static-token budget $k$ (with $\tau{=}0.5$).
+Method SR (%) $\uparrow$ FLOPs (T) $\downarrow$ Latency (ms) $\downarrow$ Freq. (Hz) $\uparrow$
+OpenVLA-OFT 97.8 3.995 78.35 65.44
+   + VLA-Cache ($\tau{=}0.2$) 95.6 3.384 66.93 76.74
+   + VLA-Cache ($\tau{=}0.3$) 96.2 3.283 67.03 79.31
+   + VLA-Cache ($\tau{=}0.4$) 98.0 3.204 66.38 79.79
+   + VLA-Cache ($\tau{=}0.5$) 98.2 3.156 64.88 79.88
+   + VLA-Cache ($\tau{=}0.6$) 98.6 3.131 64.27 81.98
+   + VLA-Cache ($\tau{=}0.7$) 98.4 3.068 63.12 82.96
+Table 10: Varying the relevance threshold $\tau$ (with k=100k{=}100).
+Overall, efficiency (FLOPs and latency) improves monotonically with larger $k$ and $\tau$, while success rate remains consistently high, corroborating the stability of VLA-Cache across sensitivity settings.
+Our default $(k{=}100,\ \tau{=}0.5)$ is used for all main results unless stated otherwise.
+Applicability to diffusion-based and alternative VLA architectures.
+VLA-Cache is designed to accelerate the language-decoder stage of VLAs and is therefore not directly applicable to models without a vision–language backbone, such as standalone diffusion policies.
+However, for hybrid architectures that combine a VLM with a diffusion-based policy head (e.g., CogACT [19]), our method remains fully compatible and has shown consistent efficiency gains and stable task performance in the SIMPLER environment, as shown in Table 3.
+Since VLA-Cache operates by reusing temporally static visual tokens before action generation, it can be integrated with diffusion-based or transformer-style models (e.g., RDT, DiT) that exhibit inter-frame redundancy, offering a promising direction for future generalization.
+E.4 Real Robot Experiment
 Robot Setup.
@@
 Data Preprocessing.
-The preprocessing procedure was conducted as follows: Initially, the recorded frames underwent center cropping, reducing the resolution from 1280×720 to 912×720, followed by resizing to 224×224.
-Subsequently, episodes were manually selected based on a visual inspection of the video sequences generated from the recorded frames.
-To minimize potential biases associated with excessively long episodes, those exceeding 250 steps were excluded from the dataset.
-Furthermore, steps in which all recorded action values were zero were removed to ensure data relevance and integrity.
-A.4 Simulated Evaluation Detail.
-LIBERO Task Definitions.
-Similarly, we also utilize all task suites provided in LIBERO for our evaluations.
-The Robosuite-based robot setup includes the following tasks: 1) “place bowl on plate with spatial variation” (e.g., drawer positions), 2) “pick object” (e.g., ketchup, bowl, apple), 3) “(open / close) target drawer; action object” (e.g., “open top drawer; place apple into drawer”), and 4) “achieve goal using shared objects” (e.g., rearranging spatial relationships or altering object states).
-SIMPLER Task Definitions.
-We utilize all task variants provided in SIMPLER for our evaluations, which include the Google robot setup with the following tasks: 1) “pick Coke can”, 2) “move obj1 near obj2”, 3) “(open / close) (top / middle / bottom) drawer”, and 4) “open top drawer; place apple into top drawer”.
-Evaluations for the Google robot setup are provided for both Visual Matching (VM) and Variant Aggregations (VA).
-Implementation Details.
-Simulated evaluations for CogACT and SIMPLER are conducted on a single NVIDIA RTX 4090 GPU in BF16 precision.
-During inference, we use DDIM sampling with 10 steps and a classifier-free guidance (CFG) coefficient of 1.5.
-Similarly, for OpenVLA and LIBERO, inference is performed on a single NVIDIA RTX 4090 GPU in BF16 precision.
-A.5 Additional Simulation Results
-Result of Subtask on LIBERO Spatial Task Suit.
 […]
-task1 task2 task3 task4 task5 task6 task7 task8 task9 task10 Avg
-Baseline (OpenVLA) 90 90 84 96 70 90 96 76 82 70 84.4 1.888 52.37
-SparseVLM 88 60 90 90 60 82 90 92 72 74 79.8 1.367 88.08
-FastV 92 90 90 94 58 92 90 80 78 70 83.4 1.888 54.00
-VLA-Cache 90 90 88 94 66 84 94 84 76 72 83.8 1.382 32.22
-Table 6: Comparison of success rates across different tasks in the LIBERO-Spatial benchmark.
-Table 6 presents detailed results on each subtask in the LIBERO-Spatial suite.
-We observe that VLA-Cache, along with methods like SparseVLM and FastV, occasionally surpasses the baseline’s success rate on individual subtasks.
-This suggests that certain redundant tokens may distract the baseline model, and pruning or reusing tokens can in fact enhance its robustness.
-Figure 7: VLA-Cache test results and attention heat map in a simulated environment
-Visualization Results.
-Figure 7 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache in LIBERO.
-A.6 Real-World Evaluation Detail.
-Task Definitions.
+Task Definitions and Success Criteria.
+We design four single-instruction tasks for the real-world robot setting to evaluate manipulation performance across diverse object types and interaction modes. The task definitions are as follows:
   1. 1.
-    Pick Up Orange Pot (single-instruction): The robot's goal is to grasp the orange pot and lift it completely off the table.
+    Pick Up Orange Pot (PickPot): The robot's goal is to grasp the orange pot and lift it completely off the table.
+A trial is successful if the pot is grasped and raised with visible clearance from the surface.
 We collected 218 valid demonstrations for the training dataset, with slight random adjustments to the initial positions of the pot and the robotic arm in each episode.
@@
   2. 2.
-    Place Blue Cube in Box (single-instruction): The robot's goal is to place the held blue cube into the target container.
+    Place Blue Cube in Box (PlaceCube): The robot's goal is to place the held blue cube into the target container.
+A trial is successful if the cube is fully inside the container at the terminal step without contact-induced ejection.
 We collected 212 valid demonstrations for the training dataset, with randomized container positions and robotic arm initial configurations in each episode.
@@
   3. 3.
-    Put Sausage in Blue Pan (single-instruction): The robot's goal is to stably place the held sausage into the blue pan.
+    Put Sausage in Blue Pan (PutSausage): The robot's goal is to stably place the held sausage into the blue pan.
+Success requires the sausage to rest inside the pan boundary at the terminal step.
 We collected 219 valid demonstrations for the training dataset, with randomized pan coordinates and robotic arm joint angles across trials.
@@
   4. 4.
-    Wipe Table (single-instruction): The robotic arm's goal is to sweep scattered items into a fixed-position dustpan using a broom.
+    Wipe Table (WipeTable): The robotic arm's goal is to sweep scattered items into a fixed-position dustpan using a broom.
+Success requires all designated items to be inside the dustpan area at termination.
 For the training dataset, we collected 187 valid demonstrations featuring randomized placements of simulated items (e.g., fries/cheese) on the table and varied initial joint configurations of the robotic arm, while the dustpan location remained fixed.
@@
 During evaluation, each trial is recorded as a success (1) or failure (0); there is no partial credit.
+Trial Protocol and Randomization.
+We run 20 trials for PickPot and PutSausage, and 30 trials for PlaceCube and WipeTable, for a total of 100 trials per method.
+For each trial we randomize the initial robot configuration and the object placement within a bounded region of the workspace.
+Trials are single-attempt with a fixed horizon; there is no human intervention or reset within a trial.
+Outcomes are recorded as binary success or failure per the criteria above.
+OpenVLA OpenVLA + VLA-Cache
+Task Success Failure SR (%) Success Failure SR (%)
+PickPot (20) 19 1 95.0 18 2 90.0
+PlaceCube (30) 25 5 83.3 27 3 90.0
+PutSausage (20) 16 4 80.0 17 3 85.0
+WipeTable (30) 21 9 70.0 22 8 73.3
+Total (100) 81 19 82.1 84 16 84.6
+Table 11: Real-world results with trial counts and success rates.
+Results with Counts and Rates.
+Table 11 reports per-task successes and failures, along with the corresponding success rate. Average success is computed across all $100$ trials.
+Figure 7: Comparison between baseline (OpenVLA) and VLA-Cache in real environment.
 Visualization Results.
-Figure 8 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache on the Kinova Jaco Robot Arm.
-Figure 8: Comparison between baseline(OpenVLA) and VLA-Cache in real environment
+Figure 7 present evaluation examples of each tasks executed by OpenVLA with VLA-Cache on the Kinova Jaco Robot Arm. More visualized results are available in the supplementary materials.